### v1

# Surrogate to Poincaré inequalities on manifolds for dimension reduction in nonlinear feature spaces

Abstract
We aim to approximate a continuously differentiable function $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ by a composition of functions $f\circ g$ where $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$, $m\leq d$, and $f:\mathbb{R}^{m}\rightarrow\mathbb{R}$ are built in a two stage procedure. For a fixed $g$, we build $f$ using classical regression methods, involving evaluations of $u$. Recent works proposed to build a nonlinear $g$ by minimizing a loss function $\mathcal{J}(g)$ derived from Poincaré inequalities on manifolds, involving evaluations of the gradient of $u$. A problem is that minimizing $\mathcal{J}$ may be a challenging task. Hence in this work, we introduce new convex surrogates to $\mathcal{J}$. Leveraging concentration inequalities, we provide sub-optimality results for a class of functions $g$, including polynomials, and a wide class of input probability measures. We investigate performances on different benchmarks for various training sample sizes. We show that our approach outperforms standard iterative methods for minimizing the training Poincaré inequality based loss, often resulting in better approximation errors, especially for rather small training sets and $m=1$.
Keywords.
high-dimensional approximation, Poincaré inequality, dimension reduction, nonlinear feature learning, deviation inequalities.
MSC Classification.
65D40, 65D15, 41A10, 41A63, 60F10
1 Introduction
We consider the problem of approximating a function $u:\mathbb{R}^{d}\to\mathbb{R}$ by a composition $f\circ g$ with $g:\mathbb{R}^{d}\to\mathbb{R}^{m}$ a feature map which associates to $\mathbf{x}\in\mathbb{R}^{d}$ a set of $m$ real features $g(\mathbf{x})=(g_{1}(\mathbf{x}),\ldots,g_{m}(\mathbf{x}))\in\mathbb{R}^{m}$. The map $g$ is selected from some model class $\mathcal{G}_{m}$. We assume that $u$ is in $L^{2}(\mathbb{R}^{d},\mu)$, where $\mu$ is some probability measure, which we equip with the natural norm $\|u\|^{2}:=\mathbb{E}(|u(\mathbf{X})|^{2})$, where $\mathbf{X}$ is a random vector with distribution $\mu$. For a fixed function $g$, an optimal $L^{2}$-approximation of the form $f\circ g$ is a solution of the optimization problem
$\min_{f}\mathbb{E}\left[|u(\mathbf{X})-f\circ g(\mathbf{X})|^{2}\right]:=\mathcal{E}(g),$ (1.1)
where the infimum is taken over all measurable maps $f$ from $\mathbb{R}^{m}$ to $\mathbb{R}$. It admits as a solution the regression function $f^{\star}(\mathbf{z})=\mathbb{E}(u(\mathbf{X})|g(\mathbf{X})=\mathbf{z})$. An optimal feature map $g$ from the class $\mathcal{G}_{m}$ can then be defined by minimizing the error $\mathcal{E}(g)$ over $\mathcal{G}_{m}$. This results in an optimal dimension reduction.
The simplest model class $\mathcal{G}_{m}$ one could think of are linear feature maps $g(\mathbf{x})=G^{T}\mathbf{x}$ for some matrix $G\in\mathbb{R}^{d\times m}$, which can be referred to as linear dimension reduction, and the associated approximation of $u$ is then a ridge function [21]. The statistical regression literature denoted this framework as (linear) sufficient dimension reduction [2], in which the main idea is to find a small $m$-dimensional subspace $\text{span}\{G\}$, spanned by the columns of $G\in\mathbb{R}^{d\times m}$, such that the random variables $u(\mathbf{X})$ and $\mathbf{X}|G^{T}\mathbf{X}$ are independent. This subspace is generally built using methods based on conditional variance, such as sliced inverse regression (SIR) [18], sliced average variance estimation (SAVE) [9] or directional regression (DR) [17]. A main issue of such methods is that they do not provide certification on the approximation error $\mathcal{E}(g)$. On the other hand, the so-called active subspace method [8, 27], which is a gradient-based method closely related to the present work, provides such a certification by leveraging probabilistic Poincaré inequalities. The latter yields an upper bound of the form
$\mathcal{E}(g)\leq C(\mathbf{X}|\mathcal{G}_{m})\mathcal{J}(g)$ (1.2)
where $\mathcal{J}(g)=\mathbb{E}\left[\|\Pi^{\perp}_{\text{span}\{G\}}\nabla u(\mathbf{X})\|_{2}^{2}\right]$ and $C(\mathbf{X}|\mathcal{G}_{m})$ is a constant depending only on the law of $\mathbf{X}$ and the model class $\mathcal{G}_{m}$. The constant $C(\mathbf{X}|\mathcal{G}_{m})$ is defined via probabilistic Poincaré constants associated to the conditional random variables $\mathbf{X}|G^{T}\mathbf{X}=\mathbf{z}$ for $\mathbf{z}\in\mathbb{R}^{m}$. The fact that the support of the latter lies in the $(d-m)$-dimensional affine space $\mathbf{z}+\text{span}\{(\}G)^{\perp}$ allows to derive bounds on $C(\mathbf{X}|\mathcal{G}_{m})$ for several classical probability measures [23]. The other main advantage of this method is that, while minimizing $\mathcal{E}$ on $\mathbb{R}^{d\times m}$ is a challenging problem, minimizing $\mathcal{J}$ on $\mathbb{R}^{d\times m}$ is rather straightforward. Indeed, it is a quadratic functional, whose minimum is the sum of the $d-m$ smallest eigenvalues of the matrix $\mathbb{E}\left[\nabla u(\mathbf{X})\nabla u(\mathbf{X})^{T}\right]\in\mathbb{R}^{d\times d}$, and the minimizer is a matrix whose columns are the eigenvectors associated with the $m$ largest eigenvalues of that matrix.
Although linear dimension reduction has several theoretical and practical advantages, the main limitation is that it may not efficiently detect important nonlinear features, as it is easy to see by considering $u()\mathbf{x})=\mapsto h(\|\mathbf{x}\|^{2}_{2})$ for some $h\in\mathcal{C}^{1}$. In this regard, there are insights for developing nonlinear dimension reduction methods, where the model class $\mathcal{G}_{m}$ contains nonlinear functions. Until recently, there were mainly nonlinear sufficient dimension reduction methods developed, see for example [26, 14, 15, 16], extending the linear sufficient dimension reduction methods (SIR, SAVE, DR). More recently, various works introduced gradient-based nonlinear dimension reduction methods, either in a heuristic manner [28, 7, 22], or using Poincaré inequalities [3, 25]. In particular in [3] it was shown that (1.2) still holds, under important assumptions, with
$\mathcal{J}(g)=\mathbb{E}\left[\|\Pi^{\perp}_{\text{span}\{\nabla\}g(\mathbf{X})}\nabla u(\mathbf{X})\|^{2}_{2}\right]=\mathbb{E}\left[\|\nabla u(\mathbf{X})\|^{2}_{2}\right]-\mathbb{E}\left[\|\Pi_{\text{span}\{\nabla\}g(\mathbf{X})}\nabla u(\mathbf{X})\|^{2}_{2}\right].$ (1.3)
The constant $C(\mathbf{X}|\mathcal{G}_{m})$ is here defined via probabilistic Poincaré inequalities on manifolds, associated to the conditional random variables $\mathbf{X}|g(\mathbf{X})=\mathbf{z}$ for $\mathbf{z}\in\mathbb{R}^{m}$. Indeed, the support of the latter lies in $g^{-1}(\{\mathbf{z}\})$, which is no longer an affine space but rather a $(d-m)$-dimensional smooth submanifold of $\mathbb{R}^{d}$, under some assumptions on $g$. Note that $\Pi^{\perp}_{\text{span}\{\nabla\}g(\mathbf{X})}\nabla u(\mathbf{X})$ is actually the Riemanian gradient of $u_{|g^{-1}(\{g(\mathbf{X})\})}$ at point $\mathbf{X}$, associated to the euclidean metric in $\mathbb{R}^{d}$.
This nonlinear setting comes with two main issues. The first is that the theoretical foundation of (1.2) is much weaker than for linear features because controlling, or simply ensuring finiteness, for $C(\mathbf{X}|\mathcal{G}_{m})$ is a truly challenging task. A first step in overcoming this problem has been proposed in [25] by considering diffeomorphisms-based feature maps $g:\mathbf{x}\mapsto(\varphi_{1}(\mathbf{x}),\cdots,\varphi_{m}(\mathbf{x}))$, with $\varphi$ a $\mathcal{C}^{1}$ diffeomorphism on $\mathbb{R}^{d}$. The second issue is that optimizing $\mathcal{J}$ over a class $\mathcal{G}_{m}$ of nonlinear functions can be a challenging task, although first and second order optimization algorithms are available. This is mainly due to the nonlinearity of the mapping $g\mapsto\Pi^{\perp}_{\text{span}\{\nabla\}g(\mathbf{x})}$. This second issue is the motivation for the present work, in which we derive new upper bounds that yield to straightforward optimization procedures.
Finally, we emphasize the fact that the approaches described in this introduction, as well as the one developed in the present paper, are two steps procedures. Firstly, we build the feature map $g$. Secondly, since the best regression function $f^{*}$ is not available in practice, we build $f:\mathbb{R}^{m}\rightarrow\mathbb{R}$ by minimizing $\mathbb{E}\left[(u(\mathbf{X})-f\circ g(\mathbf{X}))^{2}\right]$ over some regression class $\mathcal{F}_{m}$. As a result, the construction of $g$ is oblivious to the class $\mathcal{F}_{m}$, which might be problematic in practice. Another approach would be to optimize on $f$ and $g$ in a coupled way, as for example proposed in [12, 13]. We also emphasis the fact that the approaches based on Poincaré inequalities require samples of $u(\mathbf{X})$ as well as $\nabla u(\mathbf{X})$.
Contributions
The main contribution of the present work is to introduce new surrogates in order to circumvent the challenge of optimizing $\mathcal{J}$. Those surrogates are much easier to optimize, and their suboptimality with respect to $\mathcal{J}$ can be somehow controlled, assuming that $\mathcal{G}_{m}$ satisfies some deviation inequalities. We mainly provide results for the one feature case $m=1$, for which we introduce the following new surrogate,
$\mathcal{L}_{1}(g)=\mathbb{E}\left[\|\nabla u(\mathbf{X})\|_{2}^{2}\|\Pi^{\perp}_{\nabla u(\mathbf{X})}\nabla g(\mathbf{X})\|_{2}^{2}\right].$
If $\mathcal{G}_{1}$ satisfies some assumptions allowing to draw deviation inequalities on $\|\nabla g(\mathbf{X})\|_{2}^{2}$, typically polynomials, then we show that
$\mathcal{L}_{1}(g)\lesssim\mathcal{J}(g)\lesssim\mathcal{L}_{1}(g)^{\beta},$
with $\mathcal{L}_{1}$ a quadratic function and $\beta\in(0,1]$ depending on the rate of decay involved in the deviation inequalities. We also show that that if $g(\mathbf{x})=G^{T}\Phi(\mathbf{x})$ for some $\Phi\in\mathcal{C}^{1}(\mathbb{R}^{d},\mathbb{R}^{K})$ and some $G\in\mathbb{R}^{K}$ then $\mathcal{L}_{1}(g)=G^{T}HG$ with
$H=\mathbb{E}\left[\nabla\Phi(\mathbf{X})^{T}\big{(}\|\nabla u(\mathbf{X})\|_{2}^{2}I_{d}-\nabla u(\mathbf{X})\nabla u(\mathbf{X})^{T}\big{)}\nabla\Phi(\mathbf{X})\right]\in\mathbb{R}^{K\times K},$
which means that minimizing $\mathcal{L}_{1}$ is equivalent to finding an eigenvector associated to the smallest eigenvalue of $H$.
We also propose an extension of our approach to the case of multiple features $(m>1)$. We first introduce for any $1\leq j\leq m$ a new surrogate quantity $\mathcal{L}_{m,j}:\mathcal{G}_{m}\rightarrow\mathbb{R}$ which is quadratic with respect to $g_{j}$. We show that for polynomial features this surrogate satisfies similar suboptimality properties as in the case $m=1$, but with a deteriorated exponent $\beta$. We then propose a greedy algorithm to sequentially learn $g_{1},\cdots,g_{m}$, where at iteration $j$ of the algorithm we choose $g_{j}$ as a minimizer of $h\mapsto\mathcal{L}_{j,j}(g_{1},\cdots,g_{j-1},h)$.
In practice, the feature maps $g$ obtained by minimizing our surrogates for $m=1$ or $m>1$ can be used as good starting guesses of descent algorithms for minimizing $\mathcal{J}$, or directly as feature maps for building the regression function $f:\mathbb{R}\rightarrow\mathbb{R}$.
The rest of this paper is organized as follows. First in LABEL:sec:Dimension_reduction_using_Poincar\'e_inequalities we summarize the nonlinear dimension reduction method based on Poincaré inequalities on manifolds introduced in [3], and we discuss on particular choices for the model class $\mathcal{G}_{m}$. In Section 3 we shortly summarize and adapt to our framework the deviation inequalities from [10], which we leverage to ensure theoretical guarantees on our surrogates. Then in Section 4 we introduce our new surrogate $\mathcal{L}_{1}$ for $m=1$ and provide a general analysis. In Section 5 we extend the method to the case $m>1$ using a greedy algorithm. Finally in Section 6 we illustrate our methods on numerical examples.
2 Dimension reduction using Poincaré inequalities
In this section, we detail how one can leverage Poincaré inequalities to construct a feature map $g$ for the approximation of $u$ by a composition $f\circ g$, with $g$ from some model class $\mathcal{G}_{m}$ and $f$ some measurable map from $\mathbb{R}^{m}$ to $\mathbb{R}$. First in Section 2.1 we present the nonlinear dimension reduction approach proposed in [3]. Then in Section 2.2 we discuss on different choices of model classes $\mathcal{G}_{m}$.
2.1 Error bounds based on Poincaré inequalities
A key observation is that for any $g\in\mathcal{C}^{1}(\mathcal{X},\mathbb{R}^{m})$ and $f\in\mathcal{C}^{1}(\mathbb{R}^{m},\mathbb{R})$, applying the chain rule to $u=f\circ g$ yields $\nabla u(\mathbf{x})=\nabla g(\mathbf{x})\nabla f(g(\mathbf{x}))$, where $\nabla g(\mathbf{x})\in\mathbb{R}^{d\times m}$ is the jacobian matrix of $g$. Therefore,
$\nabla u(\mathbf{x})\in\text{span}\{\nabla g(\mathbf{x})\}:=\text{span}\{\nabla g_{i}(\mathbf{x})\}_{1\leq i\leq m}\subset\mathbb{R}^{d}.$
Hence there is an incentive to build $g$ such that $\nabla u(\mathbf{x})$ and $\text{span}\{\nabla\}g(\mathbf{x})$ are as aligned as possible. The authors in [3] propose to minimize the following cost function,
$\mathcal{J}(g):=\mathbb{E}\left[\|\nabla u(\mathbf{X})\|^{2}\right]-\mathbb{E}\left[\|\Pi_{\nabla g(\mathbf{X})}\nabla u(\mathbf{X})\|^{2}\right],$ (2.1)
where $\Pi_{\nabla g(\mathbf{x})}:=\Pi_{\text{span}\{\nabla\}g(\mathbf{x})}\in\mathbb{R}^{d\times d}$ denotes the orthogonal projector from $\mathbb{R}^{d}$ onto $\text{span}\{\nabla\}g(\mathbf{x})$. The cost function $\mathcal{J}(g)$ naturally appears when using Poincaré inequalities and provides certified error bounds for $g\in\mathcal{G}_{m}$ [27]. When $\mathcal{G}_{m}$ contains nonlinear functions, these error bounds involve Poincaré inequalities on Riemannian manifolds, which we define in Definition 2.1 as in [3, 25]. Deriving these inequalities involve concepts from differential geometry, and although a deep understanding of the latter is not required to read the current section, we recommend [1, Chapter 3] for a comprehensive introduction to differential geometry and its applications.
Definition 2.1 (Poincaré inequality on Riemannian submanifold of $\mathbb{R}^{d}$).
Let $\mathbf{Y}$ a continuous random vector taking values in a Riemannian submanifold $\mathcal{M}\subset\mathbb{R}^{d}$ equipped with the euclidean metric. Then the Poincaré constant $C(\mathbf{Y})\geq 0$ is the smallest constant such that
$\mathbb{E}\left[(h(\mathbf{Y})-\mathbb{E}\left[h(\mathbf{Y})\right])^{2}\right]\leq C(\mathbf{Y})\mathbb{E}\left[\|\nabla h(\mathbf{Y})\|_{2}^{2}\right]$ (2.2)
holds for any $h\in\mathcal{C}^{1}(\mathcal{M},\mathbb{R})$. Here $\nabla h$ denotes the Riemannian gradient. $\mathbf{Y}$ is said to satisfy Poincaré inequality if $C(\mathbf{Y})<+\infty$.
One of the differences between (2.2) and the classical Poincaré inequality in euclidean spaces is that it involves the Riemannian gradient $\nabla h(\mathbf{y})\in T_{\mathbf{y}}\mathcal{M}$, where $T_{\mathbf{y}}\mathcal{M}$ is the tangent space of $\mathcal{M}$ at $\mathbf{y}\in\mathcal{M}$. Fortunately, the fact that $\mathcal{M}$ is embedded in $\mathbb{R}^{d}$ simplifies some concepts involved. Indeed, we can identify $T_{\mathbf{y}}\mathcal{M}$ with a linear subspace of $\mathbb{R}^{d}$ and express the Riemannian gradient of $h\in\mathcal{C}^{1}(\mathcal{M},\mathbb{R})$ as
$\nabla h(\mathbf{y})=\Pi_{T_{\mathbf{y}}\mathcal{M}}\nabla\bar{h}(\mathbf{y}),$
where $\Pi_{T_{\mathbf{y}}\mathcal{M}}$ is the orthogonal projection onto $T_{\mathbf{y}}\mathcal{M}$, and $\bar{h}\in\mathcal{C}^{1}(\mathbb{R}^{d},\mathbb{R})$ is a function whose restriction $h_{|\mathcal{M}}$ to $\mathcal{M}$ is equal to $h$.
In our setting, the manifolds involved are the level-sets of $g$, which are smooth manifolds of dimensions $d-m$, assuming that $\text{rank}(\nabla g(\mathbf{x}))=m$ for all $\mathbf{x}\in\mathcal{X}$. The manifold-valued random vector involved is the conditional random variable $\mathbf{Y}=(\mathbf{X}|g(\mathbf{X})=\mathbf{z})$ for some $\mathbf{z}\in g(\mathcal{X})$, which takes values in the manifold $\mathcal{M}_{\mathbf{z}}=g^{-1}(\{\mathbf{z}\})$ (a level-set of $g$), and the function defined on $\mathcal{M}_{\mathbf{z}}$ is $u_{|\mathcal{M}_{\mathbf{z}}}$. In this case, for any $\mathbf{x}\in\mathcal{M}_{\mathbf{z}}$, it holds
$T_{x}\mathcal{M}_{\mathbf{z}}=\text{ker}\nabla g(\mathbf{x})^{T}=(\mathrm{span}\nabla g(\mathbf{x}))^{\perp}$
and thus
$\nabla(u_{|\mathcal{M}_{\mathbf{z}}})(\mathbf{x})=\nabla u(\mathbf{x})-\Pi_{\nabla g(\mathbf{x})}\nabla u(\mathbf{x}).$
Now applying (2.2) to $u_{|\mathcal{M}_{\mathbf{z}}}$ and taking the supremum over $\mathbf{z}\in g(\mathcal{X})$ and $g\in\mathcal{G}_{m}$ of the Poincaré constant $C(\mathbf{X}|g(\mathbf{X})=\mathbf{z})$ yields Proposition 2.1, which is a restatement of [3, Proposition 2.9].
Proposition 2.1.
Let $\mathcal{G}_{m}\subset\mathcal{C}^{1}(\mathcal{X},\mathbb{R}^{m})$ satisfying rank$(\nabla g(\mathbf{x}))=m$ for all $g\in\mathcal{G}_{m}$ and all $\mathbf{x}\in\mathcal{X}$, and assume that
$C(\mathbf{X}|\mathcal{G}_{m}):=\sup_{g\in\mathcal{G}_{m}}\sup_{z\in g(\mathcal{X})}C(\mathbf{X}|g(\mathbf{X})=\mathbf{z})<+\infty,$ (2.3)
with $C(\mathbf{X}|g(\mathbf{X})=\mathbf{z})$ defined in (2.2). Then for any $u\in\mathcal{C}^{1}(\mathcal{X},\mathbb{R})$ and $g\in\mathcal{G}_{m}$ it holds
$\mathcal{E}(g)\leq C(\mathbf{X}|\mathcal{G}_{m})\mathcal{J}(g),$ (2.4)
with $\mathcal{E}$ defined in (1.1) and $\mathcal{J}$ defined in (2.1).
The bound from Proposition 2.1 is the theoretical justification behind the idea of minimizing $\mathcal{J}$ in order to build the nonlinear feature map $g$. However, this strategy might be far from optimal since we are minimizing an upper bound of the true reconstruction error $\mathcal{E}$, especially if the constant $C(\mathbf{X}|\mathcal{G}_{m})$ is large. Still, the numerical results in [3, 25] show good performances of such Poincaré inequalities based approach on various practical examples.
It is important to note that $\mathcal{J}$ satisfies some invariance properties. Indeed, for any diffeomorphism $\psi:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$, by the chain rule and invertibility of $\nabla\psi(\mathbf{x})$, it holds that $\text{span}\{\nabla\}(\psi\circ g(\mathbf{x}))=\text{span}\{\nabla\}g(\mathbf{x})$ and thus
$\mathcal{J}(\psi\circ g)=\mathcal{J}(g).$ (2.5)
This defines an equivalence relation on $\mathcal{G}_{m}$, which allows in practice to restrict the set $\mathcal{G}_{m}$. In particular, since $\mathcal{J}(g/\|g\|)=\mathcal{J}(g)$ for any $g\neq 0$, one can impose that $\mathcal{G}_{m}$ is contained in the unit sphere of some normed space. Also, up to a suitable affine transformation $\psi$, one can consider that $\mathbb{E}\left[g(\mathbf{X})\right]=0$ and $\mathbb{E}\left[g(\mathbf{X})g(\mathbf{X})^{T}\right]=I_{m}$, which is imposed by the authors in [3] for improving numerical stability of algorithms. However, note that restricting the function class $\mathcal{G}_{m}$ based on the above invariance property of $\mathcal{J}$ may have an impact when it comes to the practical construction of the function $f$ in the composition $f\circ g$.
Finally, it is worth mentioning that the method detailed in the current section can be extended naturally to vector valued functions, as detailed in [25].
2.2 Choice of feature maps
In this section we discuss classes of feature maps that have been proposed in the framework described in Section 2.1.
2.2.1 Linear feature maps
Linear features correspond to the linear space
$\mathcal{G}_{m}=\{g:\mathbf{x}\mapsto G^{T}\mathbf{x}:G\in\mathbb{R}^{d\times m}\}.$ (2.6)
It corresponds to the so-called active subspace method [8, 27]. It has been used successfully in various practical applications.
This setting has some advantages. First, there exist known bounds on $C(\mathbf{X}|\mathcal{G}_{m})$ for some standard probability distributions for $\mathbf{X}$. For example, if $\mathbf{X}\sim\mathcal{N}(0,I_{d})$ then one can show that $C(\mathbf{X}|\mathcal{G}_{m})=1$. Secondly, a minimizer of $\mathcal{J}$ can be obtained by computing the leading eigenvectors of the matrix $\mathbb{E}\left[\nabla u(\mathbf{X})\nabla u(\mathbf{X})^{T}\right]\in\mathbb{R}^{d\times d}$. However, choosing linear features restricts the class of functions $u$ that can be well approximated. As an example, the function $\mathbf{x}\mapsto\sin(\|\mathbf{x}\|_{2}^{2})$ can’t be well approximated using a few linear features.
2.2.2 Feature maps from a linear space
Feature maps can be chosen from some finite dimensional linear space
$\mathcal{G}_{m}=\{g:\mathbf{x}\mapsto G^{T}\Phi(\mathbf{x}):G\in\mathbb{R}^{K\times m}\}.$ (2.7)
Here $\Phi\in\mathcal{C}^{1}(\mathbb{R}^{d},\mathbb{R}^{K})$, with $K\geq d$, is some fixed (or adaptively selected) feature map. The coordinates $g_{i}$ of $g\in\mathcal{G}_{m}$ are in the $K$-dimensional linear space $\text{span}\{\Phi_{1}(\mathbf{x}),\ldots,\Phi_{K}(\mathbf{x})\}$. This has been considered in [3, 22]. It has the advantage of being potentially able to approximate functions with fewer features than with (2.6), while keeping a simple parameterization of $\mathcal{G}_{m}$. Moreover we can compute analytically first and second order derivatives of $G\mapsto\mathcal{J}(G^{T}\Phi)$ and exploit the invariance property (2.5) to add constrains to the minimization procedure. In particular $G\mapsto\mathcal{J}(G^{T}\Phi)$ actually only depends on $\text{span}\{G\}\subset\mathbb{R}^{K}$. Thus one can for example consider instead the compact set
$\mathcal{G}_{m}=\Big{\{}g:\mathbf{x}\mapsto G^{T}\Phi(\mathbf{x}):G\in\mathbb{R}^{K\times d},G^{T}RG=I_{m}\Big{\}},$ (2.8)
with $R\in\mathbb{R}^{K\times K}$ a symmetric positive definite matrix. The set being identified with a Stiefel matrix manifold, optimization algorithms on manifolds could be used [1]. One may go even further and restrict the optimization to the Grassmann manifold of $m$-dimensional subspaces in $\mathbb{R}^{K}$, denoted $Grass(m,K)$. Up to our knowledge, such algorithms have not been considered yet.
There are two main issues with this class of feature maps. The first is theoretical and lies in the constant $C(\mathbf{X}|\mathcal{G}_{m})$, since there is currently no way (even heuristic) to control it. Worse, if the level set $g^{-1}(\{\mathbf{z}\})$ is not smoothly pathwise-connected, then we know that $C(\mathbf{X}|g(\mathbf{X})=\mathbf{z})=+\infty$, which breaks down the theoretical guarantees of the method. This can be the case for example when using as feature maps polynomials of degree 2. The second issue is that although the coordinates of $g\in\mathcal{G}_{m}$ are linear combinations of fixed features $\Phi(\mathbf{X})$, minimizing $G\mapsto\mathcal{J}(G^{T}\Phi)$ is much more difficult than in the linear case. It does not boil down to solving an eigenvalue problem, which is mainly due to the fact that the projection matrix $\Pi_{\nabla g(\mathbf{x})}$ depends on $\mathbf{x}$. Still, despite these two issues, the numerical results in [3] give strong positive incentives for using this method, even when connectivity of level sets is not satisfied.
2.2.3 Feature maps based on diffeomorphisms
In [28, 25], the authors considered for $\mathcal{G}_{m}$ nonlinear feature maps defined as the first $m$ components of a class of $\mathcal{C}^{1}$-diffeomorphisms, i.e.
$\mathcal{G}_{m}=\{g:\mathbf{x}\mapsto(\varphi_{1}(\mathbf{x}),\cdots,\varphi_{m}(\mathbf{x})):\varphi\in\mathcal{D}\},$ (2.9)
where $\mathcal{D}$ is a set of $\mathcal{C}^{1}$-diffeomorphisms of $\mathbb{R}^{d}$, meaning that for any $\varphi\in\mathcal{D}$, it holds $\varphi,\varphi^{-1}\in\mathcal{C}^{1}(\mathbb{R}^{d},\mathbb{R}^{d})$. As with (2.7), the variety of possible feature maps $g$ has the advantage of being potentially able to approximate functions with fewer features than (2.6).
The main advantage however it that it possibly allows a better control on the Poincaré constants. In particular it ensures connectivity of the level sets of the feature maps in $\mathcal{G}_{m}$, which circumvents the main theoretical issues of (2.7) and allows to derive Poincaré inequalities for the conditional random variables $\mathbf{X}|g(\mathbf{X})=\mathbf{z}$, although the associated Poincaré constants are not available in general. Moreover, it allows to leverage Poincaré inequalities in the feature space $\varphi(\mathcal{X})$. It consists in applying Proposition 2.1 on $u\circ\varphi^{-1}:\varphi(\mathcal{X})\rightarrow\mathbb{R}$ with features $\varphi_{1}(\mathbf{X}),\cdots,\varphi_{m}(\mathbf{X})$ which are linear in $\varphi(\mathbf{X})$. As a result, if one ensures that $\varphi(\mathbf{X})\sim\mathcal{N}(0,I_{d})$, in which case $\varphi$ is called a normalizing flow for $\mathbf{X}$, then we would obtain $C(\varphi(\mathbf{X})|g(\mathbf{X})=\mathbf{z})=1$. To this end, the authors in [25] introduce a penalized optimization problem with a penalty based on some divergence (e.g., the Kullback-Leibler divergence) between $\varphi(\mathbf{X})$ and $\mathcal{N}(0,I_{d})$, which is a reasonable heuristic, although it gives no theoretical control over the Poincaré constant.
The main issue with this class of feature maps is that the set $\mathcal{G}_{m}$ has a much more complex structure than the vector space structure in (2.7). It also strongly depends on the choice of parametrization of $\mathcal{D}$. As a result, not only the objective function $\mathcal{J}$ is highly non-convex, but so is the set on which it is optimized.
We end this section by pointing out that in [25] the authors proposed a dimension augmentation strategy, consisting in approximating $u(\mathbf{x})$ by $\mathbb{E}(f(g(\mathbf{x},\mathbf{\Xi}))),$ where $\mathbf{\Xi}\in\mathbb{R}^{m}$ is some random vector, independent on $\mathbf{X}$, with connected support. Then the above approach can be used to construct a map $g:\mathbb{R}^{d+m}\to\mathbb{R}^{m}$ from the first $m$ coordinates of a diffeomorphism. This approach has two main advantages. The first is that it enlarges the set $\mathcal{D}$, thus increasing the expressivity of the admissible feature maps, while preserving the theoretical guarantees. The second is that it allows to easily define diffeomorphisms $\varphi:(\mathbf{x},\mathbf{\xi})\mapsto(h(\mathbf{x})+\mathbf{\xi},\mathbf{x})$ on $\mathbb{R}^{d+m}$ from arbitrary functions $h\in\mathcal{C}^{1}(\mathbb{R}^{d},\mathbb{R}^{m})$, thus circumventing the theoretical issue of (2.7). However, it also brings some new issues, especially the fact that the approximation of $u$ involves an expectation over $\mathbf{\xi}$, see the discussion in [25].
3 Large and small deviations inequalities
In order to obtain theoretical guarantees for our new approach described in Section 4, we require so called small and large deviations inequalities for the random variables $\|\nabla g({\bf X})\|_{2}^{2}$ for $g\in\mathcal{G}_{m}$. In other words, we require upper bound on both
$\mathbb{P}\left[\|\nabla g({\bf X})\|_{2}^{2}\leq t^{-1}\right]\quad\text{and}\quad\mathbb{P}\left[\|\nabla g({\bf X})\|_{2}^{2}\geq t\right]$
as $t\rightarrow+\infty$. To that end, we summarize in this section some results from [4, 10]. Although these works provided independently similar results, we will be interested in some specific cases studied in [10], and for this reason we will mostly rely on the latter reference.
The results of interest in our framework are sharp large and small deviation inequalities for $s$-concave probability measures with $s\in[-\infty,1/d]$. This class of probability measures has been well studied. We refer to the introductions of [5, 10] and [4, Section 3] for a short summary and basic properties. We also refer to [5, 6] for a deeper study. For the sake of simplicity, here we only consider measures admitting a density with respect to the Lebesgue measure and we introduce them in Definition 3.1 via the characterization recalled in [10, Introduction].
Definition 3.1 ($s$-concave probability measure).
Let $\mu$ a probability measure on $\mathbb{R}^{d}$ such that $d\mu(\mathbf{x})=\rho(\mathbf{x})d\mathbf{x}$. For $s\in[-\infty,1/d]$, $\mu$ is $s$-concave if and only if $\rho$ is supported on a convex set and is $\kappa$-concave with $\kappa=s/(1-sd)\in[-1/d,+\infty]$, meaning
$\rho(\lambda\mathbf{x}+(1-\lambda)\mathbf{y})\geq(\lambda\rho(\mathbf{x})^{\kappa}+(1-\lambda)\rho(\mathbf{y})^{\kappa})^{1/\kappa}$ (3.1)
for all $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$ such that $\rho(\mathbf{x})\rho(\mathbf{y})>0$ and all $\lambda\in[0,1]$. The cases $s\in\{-\infty,0,1/d\}$ are interpreted by continuity.
Let us recall some basic properties of $s$-concave probability measures $\mu$ on $\mathbb{R}^{d}$. Firstly there exists no such measure for $s>1/d$, and if $\mu$ is $s$-concave, it is also $t$-concave for all $t\leq s$. Secondly for $s\in(0,1/d]$ the support of $\mu$ is compact. Also a measure is $s$-concave with $s=1/d$ if and only if the measure is uniform. Finally $s=0$ if and only if $\mu$ is log-concave. We can also note that $s=-1$ for the Cauchy distribution.
For this class of probability measures, we will state small and large deviations inequalities for functions with bounded Chebyshev degree, using the terminology (which does not seem classical) introduced in [10]. This corresponds to measurable functions $h:\mathbb{R}^{d}\rightarrow\mathbb{R}$ such that for any straight bounded segment $J\subset\mathbb{R}^{d}$ and any measurable set $I\subset J$,
$\sup_{\mathbf{x}\in J}|h(\mathbf{x})|\leq\big{(}\frac{A_{h}|J|}{|I|}\big{)}^{k_{h}}\sup_{\mathbf{x}\in I}|h(\mathbf{x})|,$ (3.2)
for some constants $k_{h}<+\infty$ and $A_{h}<+\infty$. Note that such a function $h$ also satisfies this inequality for any $k\leq k_{h}$ and $A\leq A_{h}$. On the other hand, (3.2) can be seen as a type of Remez inequality, which is more classical, therefore we will stick with the latter phrasing. Now, although there are several known classes of functions satisfying (3.2), polynomials are of particular interest. Indeed, as stated in [10], if $h=P$ is a multivariate polynomial then $P$ satisfies (3.2) with
$k_{P}=\text{deg}(P),\quad A_{P}=4,$ (3.3)
where $\text{deg}(P)$ denotes the total degree of $P$. A direct consequence is that if $\mathcal{G}_{1}$ is a set of polynomial functions with total degree at most $\ell+1$, then for any $g\in\mathcal{G}_{1}$ we have that $h:\mathbf{x}\mapsto\|\nabla g(\mathbf{x})\|_{2}^{2}$ is a multivariate, real-valued polynomial of degree $2\ell$, which thus satisfies (3.2) with constants $k_{h}=2\ell$ and $A_{h}=4$. This is stated in the following Proposition 3.1.
Proposition 3.1.
If $\mathcal{G}_{1}$ is a set of polynomials with total degree at most $\ell+1$, then for all $g\in\mathcal{G}_{1}$ the function $h:\mathbf{x}\mapsto\|\nabla g(\mathbf{x})\|_{2}^{2}$ satisfies (3.2) with constants
$k=2\ell,\quad A=4.$ (3.4)
Now, the deviations inequalities in this framework, relying on the analysis of dilatation of the sub-level sets of $\mathbf{x}\mapsto|h(\mathbf{x})|$ provided in [10, 4], allows to link the measures of two sub-level sets of the function. Then, we can choose for example a sub-level set defined by the $p$-quantile of $|h(\mathbf{X})|$, denoted by $Q_{|h(\mathbf{X})|}(p)$. We recall that for any real random variable $Z$ the quantile function is defined by
$Q_{Z}(p):=\inf\{z\in\mathbb{R}:\mathbb{P}\left[Z\leq z\right]\geq p\},\quad\forall p\in(0,1).$ (3.5)
In particular, we will only use $Q_{|h(\mathbf{X})|}(1/2)$, the median of $|h(\mathbf{X})|$. Now, since the deviations inequalities will be leveraged for a set of functions $\mathcal{K}$, we will need uniform lower and upper bounds,
$\underline{q}_{\mathcal{K}}:=\min_{h\in\mathcal{K}}Q_{|h(\mathbf{X})|}(1/2),\quad\overline{q}_{\mathcal{K}}:=\max_{h\in\mathcal{K}}Q_{|h(\mathbf{X})|}(1/2).$ (3.6)
It turns out that, in our framework, such bounds can be obtained with mild assumptions.
Proposition 3.2.
Let $\mathbf{X}$ a random variable on $\mathbb{R}^{d}$ whose law $\mu$ has connected support. If $\mathcal{K}\subset\mathcal{C}^{0}(\mathbb{R}^{d},\mathbb{R})\cap L^{1}_{\mu}(\mathbb{R}^{d},\mathbb{R})$ is a compact set in $L^{1}_{\mu}(\mathbb{R}^{d},\mathbb{R})$ then $\overline{q}_{\mathcal{K}}<+\infty$. Additionally, if for all $h\in\mathcal{K}$ it holds $\mathbb{P}\left[h(\mathbf{X})\leq 0\right]=0$, then $\underline{q}_{\mathcal{K}}>0$.
Proof.
Let us show that $\omega_{p}:h\mapsto Q_{h(\mathbf{X})}(p)$ is continuous on $\mathcal{C}^{0}\cap L^{1}_{\mu}$. From [11, Ch 14, Prop 5] we have that for all $h\in L^{1}_{\mu}$, if $Q_{h(\mathbf{X})}$ is continuous at $p\in(0,1)$, then $\omega_{p}$ is continuous at $h$. Moreover for any $h\in\mathcal{C}^{0}$, since $\mathbf{X}$ has connected support, so does $h(\mathbf{X})$, thus $Q_{h(\mathbf{X})}$ is continuous on $(0,1)$. We thus have the desired result on $\overline{q}_{\mathcal{K}}$. Finally, the assumption that $h(\mathbf{X})$ is almost surely positive for any $h\in\mathcal{K}$ directly yields the result on $\underline{q}_{\mathcal{K}}$. ∎
We can now state in Propositions 3.3 and 3.4 respectively small and large deviations results, that hold uniformly on a compact set of functions $\mathcal{K}$. Note that those results are essentially the ones from [10]. The first difference is that we slightly weakened the bounds in order to make them easier to leverage in our analysis in further sections, while not changing the rates of decay. The second difference is that we stated the results uniformly on some compact set of functions, thus we had to ensure that the constant involved can be uniformly bounded, which is ensured by Proposition 3.2.
Proposition 3.3 (Small deviations).
Let $\mathbf{X}$ an absolutely continuous random variable on $\mathbb{R}^{d}$ whose law $\mu$ is $s$-concave with $s\in[-\infty,1/d]$. Let $\mathcal{K}\subset\mathcal{C}^{0}(\mathbb{R}^{d},\mathbb{R})\cap L^{1}_{\mu}(\mathbb{R}^{d},\mathbb{R})$ a compact set in $L^{1}_{\mu}(\mathbb{R}^{d},\mathbb{R})$ such that every $h\in\mathcal{K}$ satisfies (3.2) with constants $k,A<+\infty$. Then for all $h\in\mathcal{K}$ and all $\varepsilon>0$ it holds
$\mathbb{P}\left[|h(\mathbf{X})|\leq\underline{q}\varepsilon\right]\leq\eta\varepsilon^{1/k}$ (3.7)
with $\underline{q}=\underline{q}_{\mathcal{K}}$ defined in (3.6) and $\eta=\eta^{small}_{s,\mathcal{K}}:=\max(1,(1-2^{-s})s^{-1}A)$.
Proof.
See Section A.1. ∎
Proposition 3.4 (Large deviations).
Let $\mathbf{X}$ an absolutely continuous random variable on $\mathbb{R}^{d}$ whose law $\mu$ is $s$-concave with $s\in[-\infty,1/d]$. Let $\mathcal{K}\subset\mathcal{C}^{0}(\mathbb{R}^{d},\mathbb{R})\cap L^{1}_{\mu}(\mathbb{R}^{d},\mathbb{R})$ a compact set in $L^{1}_{\mu}(\mathbb{R}^{d},\mathbb{R})$ such that every element in $\mathcal{K}$ satisfies (3.2) with constants $k,A<+\infty$. Then for all $h\in\mathcal{K}$ and all $t>0$ it holds
$\mathbb{P}\left[|h(\mathbf{X})|\geq\overline{q}t\right]\leq\left\{\begin{aligned} &\big{(}1-\frac{t^{1/k}-1}{\eta}\big{)}_{+}^{1/s},&&s\in(0,1/d]&\\ &\exp(-\frac{t^{1/k}-1}{\eta}),&&s=0,&\\ &\eta t^{1/sk},&&s\in[-\infty,0),&\end{aligned}\right.$ (3.8)
with $\overline{q}=\overline{q}_{\mathcal{K}}$ defined in (3.6) and
$\eta=\eta^{large}_{s,\mathcal{K}}:=\left\{\begin{aligned} &A/(1-2^{-s})&&s\in(0,1/d]&\\ &A/\log(2)&&s=0,&\\ &\max(1,A/(2^{-s}-1))^{-1/s}&&s\in[-\infty,0).&\end{aligned}\right.$ (3.9)
Proof.
See Section A.2
Let us end this section by considering an example. If $\mathbf{X}$ is uniformly distributed on a convex set $\mathcal{X}\in\mathbb{R}^{d}$, meaning its distribution is $s$-concave with $s=1/d$, and if $\mathcal{K}$ is a compact set of polynomials of total degree at most $\ell$, then the constants in Propositions 3.3 and 3.4 are $k=\ell$ and
$\eta^{small}_{\frac{1}{d},\mathcal{K}}=4d(1-2^{-1/d})\leq 4\log(2),\quad\eta^{large}_{\frac{1}{d},\mathcal{K}}=4/(1-2^{-1/d}).$
Also, we know that $\overline{q}_{\mathcal{K}}<+\infty$, and additionally if the polynomials in $\mathcal{K}$ are almost surely positive then from Proposition 3.2 we also know that $\underline{q}_{\mathcal{K}}>0$.
4 The case of a single feature
In this section we study the case $m=1$ where we search for a scalar-valued feature map $g$. We introduce a new quantity
$\mathcal{L}_{1}(g):=\mathbb{E}\left[\|\nabla u(\mathbf{X})\|_{2}^{2}\,\|\Pi^{\perp}_{\nabla u(\mathbf{X})}\nabla g(\mathbf{X})\|_{2}^{2}\right],\quad\forall g\in\mathcal{G}_{1},$ (4.1)
whose minimizer shall be a candidate feature map. In order for $\mathcal{L}_{1}(g)$ to be well defined, we need a first assumption on the set $\mathcal{G}_{1}$.
Assumption 4.1.
$\mathcal{G}_{1}$ is a compact subset of $\mathcal{C}^{1}(\mathbb{R}^{d},\mathbb{R})\cap W^{1,p_{1}}_{\mu}(\mathbb{R}^{d},\mathbb{R})$ for some $p_{1}>2$, endowed with the norm
$\|g\|_{p_{1}}=(\|g(\mathbf{X})\|_{p_{1}}^{p_{1}}+\|\|\nabla g(\mathbf{X})\|_{2}^{2}\|_{p_{1}}^{p_{1}})^{1/p_{1}}.$
The compactness ensures that $\mathcal{L}_{1}$ admits minimizers on $\mathcal{G}_{1}$, and that large deviations inequalities hold uniformly on $\mathcal{G}_{1}$. That being said, the idea behind considering this very surrogate lies in the following Lemma 4.2.
Lemma 4.2.
For all $\mathbf{v},\mathbf{w}\in\mathbb{R}^{d}$, it holds
$\|\mathbf{w}\|^{2}_{2}\|\Pi_{\mathbf{w}^{\perp}}\mathbf{v}\|^{2}_{2}=\|\mathbf{v}\|^{2}_{2}\|\Pi_{\mathbf{v}^{\perp}}\mathbf{w}\|^{2}_{2}.$ (4.2)
Proof.
It holds
$\|\mathbf{w}\|^{2}_{2}\|\Pi^{\perp}_{\mathbf{w}}\mathbf{v}\|^{2}_{2}=\|\mathbf{w}\|^{2}_{2}\|\mathbf{v}\|^{2}_{2}-\|\mathbf{w}\|^{2}_{2}\langle\frac{\mathbf{w}}{\|\mathbf{w}\|_{2}},\mathbf{v}\rangle^{2}_{2}=\|\mathbf{w}\|^{2}_{2}\|\mathbf{v}\|^{2}_{2}-\langle\mathbf{w},\mathbf{v}\rangle^{2}_{2}=\|\mathbf{v}\|^{2}_{2}\|\Pi^{\perp}_{\mathbf{v}}\mathbf{w}\|^{2}_{2}.$
Thus for any $g\in\mathcal{G}_{1}$, taking $\mathbf{w}=\nabla g(\mathbf{X})$ and $\mathbf{v}=\nabla u(\mathbf{X})$ in the above lemma yields
$\|\nabla g(\mathbf{X})\|^{2}\|\Pi^{\perp}_{\nabla g(\mathbf{X})}\nabla u(\mathbf{X})\|^{2}=\|\nabla u(\mathbf{X})\|^{2}\|\Pi^{\perp}_{\nabla u(\mathbf{X})}\nabla g(\mathbf{X})\|^{2}.$ (4.3)
In the right-hand side we recognize the term within the expectation in $\mathcal{L}_{1}$. In the left-hand side we recognize the term within the expectation in $\mathcal{J}$, defined by (2.1), up to the factor $\|\nabla g(\mathbf{X})\|_{2}^{2}$. As a result, controlling $\|\nabla g(\mathbf{X})\|_{2}^{2}$ is the key point for controlling the suboptimality of $\mathcal{L}_{1}$. For example for linear features $g(\mathbf{x})=G^{T}\mathbf{x}$ with $G^{T}G=1$ we have that $\|\nabla g(\mathbf{X})\|_{2}=1$ almost surely thus $\mathcal{L}_{1}(g)=\mathcal{J}(g)$. A first element of control is given by Assumption 4.3, which is formulated for any $m\geq 1$, and which implies that $\|\nabla g(\mathbf{X})\|_{2}^{2}$ is almost surely positive in the case $m=1$.
Assumption 4.3.
For all $g\in\mathcal{G}_{m}$ it holds rank$(\nabla g(\mathbf{X}))=m$ almost surely.
Finally, in order to use the large and small deviations inequalities from Section 3, we need Assumption 4.4 on $\mathcal{G}_{1}$, or more precisely on the set of real-valued functions on which we apply those inequalities, denoted by
$\mathcal{K}_{1}:=\{\mathbf{x}\mapsto\|\nabla g(\mathbf{x})\|_{2}^{2}\,:\,g\in\mathcal{G}_{1}\}.$ (4.4)
Note that by Assumption 4.1, $\mathcal{K}_{1}$ is a compact subset of $\mathcal{C}^{0}(\mathbb{R}^{d},\mathbb{R})\cap L^{p_{1}}_{\mu}(\mathbb{R}^{d},\mathbb{R})$, and by Assumption 4.3 its elements are almost surely positive.
Assumption 4.4.
For all $g\in\mathcal{G}_{m}$, the function $\mathbf{x}\mapsto\|\nabla g(\mathbf{x})\|_{2}^{2}$ satisfies (3.2) with constants $k<+\infty$ and $A<+\infty$.
One shall keep in mind that, as stated in Proposition 3.1 and the discussion after, if $\mathcal{G}_{1}$ contains only polynomial feature maps of total degree at most $\ell+1$, then $\mathcal{G}_{1}$ satisfies the above Assumption 4.4 with $k=2\ell$ and $A=4$. Additionally if $\mathcal{G}_{1}$ does not contain any constant polynomial, then it also satisfies Assumption 4.3.
The structure of the section is as follows. First in Section 4.1 we provide results on the suboptimality of $\mathcal{L}_{1}$. Then in Section 4.2 we discuss on the minimization of $\mathcal{L}_{1}$.
4.1 Suboptimality
In this section, we discuss on the suboptimality of our surrogate $\mathcal{L}_{1}$ defined in (4.1) regarding the Poincaré inequality error bound $\mathcal{J}$ defined in (2.1).
Let us start by investigating the special case when $u$ can be described with exactly one feature in $\mathcal{G}_{1}$, in other words when $\mathcal{J}(g)=0$ for some $g\in\mathcal{G}_{1}$. In this case, we can show that minimizing $\mathcal{L}_{1}$ is equivalent to minimizing $\mathcal{J}$, as stated in Proposition 4.5. Note that this result relies on the fact that there is no more than $1$ feature to learn.
Proposition 4.5.
Assume $\mathcal{G}_{1}$ satisfies Assumption 4.3. Then for all $g\in\mathcal{G}_{1}$ we have $\mathcal{J}(g)=0$ if and only if $\mathcal{L}_{1}(g)=0$.
Proof.
See Section B.1. ∎
Now in more general cases we cannot expect that $\min_{g\in\mathcal{G}_{1}}\mathcal{J}(g)=0$, and we would want to quantify the suboptimality of $\mathcal{L}_{1}$. An ideal setting, considered in the following Proposition 4.6, is when $g$ is bi-Lipschitz with constants uniformly controlled on $\mathcal{G}_{1}$. Note that this is for example the case when $\mathcal{G}_{1}$ is defined by (2.9) with $\mathcal{D}$ containing only bi-Lipschitz diffeomorphisms whose associated constants are uniformly bounded over $\mathcal{D}$.
Proposition 4.6.
Assume that there exist $c>0$ and $C<+\infty$ such that for all $g\in\mathcal{G}_{1}$ and $\mathbf{x}\in\mathcal{X}$ it holds $c\leq\|\nabla g(\mathbf{x})\|_{2}^{2}\leq C$. Then we have
$c\mathcal{J}(g)\leq\mathcal{L}_{1}(g)\leq C\mathcal{J}(g),$ (4.5)
and thus if $g^{*}$ is a minimizer of $\mathcal{L}_{1}$ over $\mathcal{G}_{1}$, it yields
$\mathcal{J}(g^{*})\leq(C/c)\inf_{g\in\mathcal{G}_{1}}\mathcal{J}(g).$ (4.6)
Proof.
The result follows directly from (4.3). ∎
In the bi-Lipschitz setting, $\mathcal{J}$ and $\mathcal{L}_{1}$ are equivalent in some sense. However this setting is very restrictive, especially regarding the lower bound on the gradient. For example if $\mathcal{G}_{m}$ is taken as (2.8) with $\Phi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{K}$ and $K>d$, then for fixed $\mathbf{x}\in\mathcal{X}$ we have rank$(\nabla\Phi(\mathbf{x}))\leq d$, thus there exists $G\in\mathbb{R}^{K}\setminus\{0\}$ such that $G^{T}\nabla\Phi(\mathbf{x})=0=\nabla g(\mathbf{x})$ where $g=G^{T}\Phi$. Thus in this case, there cannot exist bi-Lipschitz constants uniformly bounded on $\mathcal{G}_{m}$. For this reason, the theoretical guarantees on $\mathcal{L}_{1}$ will heavily rely in general on concentration-type inequalities in order to upper bound the probabilities of $(\|\nabla g(\mathbf{X})\|_{2}^{2}\geq\beta)$ and $(\|\nabla g(\mathbf{X})\|_{2}^{2}\leq\beta^{-1})$ as $\beta\rightarrow+\infty$. The first one is referred to as large deviation probability and widely known concentration inequalities (e.g., Chernoff’s) provide the desired bounds. The second one is referred to as small deviation probability, which is much less studied than large deviations.
We now introduce Sobolev-type semi-norms that will appear in our analysis of the suboptimality of $\mathcal{L}_{1}$. For a set of functions $\mathcal{A}\subset W^{1,r}(\mathbb{R}^{d},\mathbb{R})$, $r\geq 1,$ we define
$\nu_{\mathcal{A},r}:=\sup_{h\in\mathcal{A}}\nu_{h,r},\quad\nu_{h,r}:=\|\|\nabla h(\mathbf{X})\|_{2}^{2}\|_{r},$ (4.7)
with $\|\cdot\|_{r}$ the standard $L^{r}$-norm for random variables. Let us first consider the control of $\mathcal{J}$ by $\mathcal{L}_{1}$ uniformly on $\mathcal{G}_{1}$. We state such a result in Proposition 4.7, and we recall that from Proposition 3.1, if $\mathcal{G}_{1}$ contains only polynomials of total degree at most $\ell+1$ then $\mathcal{G}_{1}$ satisfies Assumption 4.4 with $k=2\ell$ and $A=4$.
Proposition 4.7.
Assume $\mathbf{X}$ is an absolutely continuous random variable on $\mathbb{R}^{d}$ whose distribution is $s$-concave with $s\in[-\infty,1/d]$. Assume that $\mathcal{G}_{1}$ satisfies Assumptions 4.3 and 4.4 and that $\nu_{u,p_{2}}<+\infty$ for some $p_{2}\geq 1$. Then denoting $p^{-1}=1-p_{2}^{-1}$ it holds
$\mathcal{J}(g)\leq\gamma_{1}\mathcal{L}_{1}(g)^{\frac{1}{1+pk}},\quad\forall g\in\mathcal{G},$ (4.8)
with $\gamma_{1}$ defined in (B.3).
Proof.
See Section B.2. ∎
We can first note that if $s>0$ then $\mathbf{X}$ has compact support, thus $\nu_{u,\infty}<+\infty$ and we can apply Proposition 4.7 with $p_{2}=\infty$ and $p=1$. Also, it is important to note that the bound (4.8) depends rather weakly on the parameter $s$ of the distribution of $\mathbf{X}$.
On the other hand, it heavily depends on the constant $k$. Now we consider the control of $\mathcal{L}_{1}$ by $\mathcal{J}$ uniformly on $\mathcal{G}_{1}$ in Proposition 4.8.
Proposition 4.8.
Assume $\mathbf{X}$ is an absolutely continuous random variable on $\mathbb{R}^{d}$ whose distribution is $s$-concave with $s\in[-\infty,1/d]$. Assume that $\mathcal{G}_{1}$ satisfies Assumptions 4.4, 4.1 and 4.3 and that $\nu_{u,p_{2}}<+\infty$ for some $p_{2}\geq 2$. Then denoting $r^{-1}=1-p_{1}^{-1}-p_{2}^{-1}$ it holds
$\mathcal{L}_{1}(g)\leq\gamma_{2}\left\{\begin{aligned} &\mathcal{J}(g)\quad&\text{if }s\in(0,1/d],\\ &\mathcal{J}(g)(1-\log(\frac{\mathcal{J}(g)}{\nu_{u,p_{2}}}))^{k}\quad&\text{if }s=0,\\ &\mathcal{J}(g)^{\frac{1}{1-srk}}\quad&\text{if }s<0,\\ \end{aligned}\right.$ (4.9)
with $\gamma_{2}$ defined in (B.4).
Proof.
See Section B.3. ∎
It is worth noting that the resulting bound strongly depends on the parameter $s$ of the law of $\mathbf{X}$, contrarily to the converse statement from Proposition 4.7. Let us briefly discuss on the different cases for $s$. First, the case $s\in(0,1/d]$ yields the best bound, and depends on $k$ only via the constant $\gamma_{2}$. This could be expected since in that case $\mathbf{X}$ has compact support, thus compactness of $\mathcal{G}_{1}$ yields $\nu_{\mathcal{G}_{1},\infty}<+\infty$. Then, the case $s=0$, which we recall is equivalent to log-concave distributions, yields a slightly worse bound due to the logarithmic term raised to the power $k$, which can still be considered as a rather weak dependency on $k$ overall. Finally, the case $s<0$ yields the worse bound, whose dependency on $k$ is roughly the same as in Equation 4.8, up to a factor $s$.
Now, we provide in Proposition 4.9 a control on the suboptimality of $\mathcal{L}_{1}$.
Proposition 4.9.
Assume $\mathbf{X}$ is an absolutely continuous random variable on $\mathbb{R}^{d}$ whose distribution is $s$-concave with $s\in[-\infty,1/d]$. Assume that $\mathcal{G}_{1}$ satisfies Assumptions 4.4, 4.1 and 4.3 and that $\nu_{u,p_{2}}<+\infty$ for some $p_{2}\geq 2$. Then denoting $p^{-1}=1-p_{1}^{-1}$ and $r^{-1}=p^{-1}-p_{2}^{-1}$ it holds
$\mathcal{J}(g^{*})\leq\gamma_{3}\left\{\begin{aligned} &\inf_{g\in\mathcal{G}_{1}}\mathcal{J}(g)^{\frac{1}{1+pk}}\quad&\text{if }s\in(0,1/d],\\ &\inf_{g\in\mathcal{G}_{1}}\mathcal{J}(g)^{\frac{1}{1+pk}}(1-\log(\frac{\mathcal{J}(g)}{\nu_{u,p_{2}}}))^{\frac{k}{1+pk}}\quad&\text{if }s=0,\\ &\inf_{g\in\mathcal{G}_{1}}\mathcal{J}(g)^{\frac{1}{1+pk}\frac{1}{1-srk}}\quad&\text{if }s<0,\\ \end{aligned}\right.$ (4.10)
where $\gamma_{3}:=\gamma_{1}\gamma_{2}^{\frac{1}{1+pk}}$ with $\gamma_{1}$ and $\gamma_{2}$ defined respectively in Propositions 4.7 and 4.8.
Proof.
Let $g^{*}\in\mathcal{G}_{1}$ be a minimizer of $\mathcal{L}_{1}$ over $\mathcal{G}_{1}$. Then for all $g\in\mathcal{G}_{1}$ we have by Proposition 4.7 that $\mathcal{J}(g^{*})\leq\gamma_{1}\mathcal{L}_{1}(g^{*})^{\frac{1}{1+pk}}\leq\gamma_{1}\mathcal{L}_{1}(g)^{\frac{1}{1+pk}}$. Now, using Proposition 4.8 and taking the infimum over $g\in\mathcal{G}_{1}$ yields the desired result. ∎
Finally, let us put a stress on the special case of polynomial feature maps with $s=1/d$, meaning when $\mathbf{X}$ is uniformly distributed on a compact convex set of $\mathbb{R}^{d}$. In this case, Proposition 4.9 becomes the much simpler Corollary 4.10 stated below.
Corollary 4.10.
Assume that $\mathbf{X}$ is uniformly distributed on a compact convex subset of $\mathbb{R}^{d}$. Assume that $\mathcal{G}_{1}$ satisfies Assumption 4.1 and contains only non-constant polynomials of total degree at most $\ell+1$. Then it holds
$\mathcal{J}(g^{*})\leq\gamma_{3}\inf_{\mathcal{G}_{1}}\mathcal{J}^{\frac{1}{1+2\ell}}$ (4.11)
with
$\gamma_{3}\leq 40d(\overline{q}/\underline{q})^{\frac{1}{1+2\ell}}\nu_{u,\infty}^{\frac{2\ell}{1+2\ell}},$ (4.12)
and $\underline{q}=\underline{q}_{\mathcal{K}_{1}}$, $\overline{q}=\overline{q}_{\mathcal{K}_{1}}$ defined in (3.6), and $\nu_{u,\infty}$ defined in (4.7).
Proof.
Since $\mathbf{X}$ is uniformly distributed on a compact convex subset of $\mathbb{R}^{d}$, its law is $s$-concave with $s=1/d$. Also, since $\mathcal{G}_{1}$ contains only non-constant polynomials of total degree at most $\ell+1$, we have that $\mathbf{x}\mapsto\|\nabla g(\mathbf{x})\|_{2}^{2}$ is a non zero polynomial of degree at most $2\ell$. As a result, $\mathcal{G}_{1}$ satisfies Assumption 4.4 with constants $k=2\ell$ and $A=4$, as well as Assumption 4.3. Moreover, since $\mathbf{X}$ has compact support, we have $\nu_{u,\infty},\nu_{\mathcal{G}_{1},\infty}<+\infty$, thus we can apply Proposition 4.9 with $s=1/d$, $p_{1}=p_{2}=+\infty$ and $p=1$. The constants $\gamma_{1},\gamma_{2}$ defined in (B.3) and (B.4) then read
$\gamma_{1}=2(\nu_{u,\infty}4d(1-2^{-1/d})\underline{q}^{-1/2\ell})^{\frac{2\ell}{1+2\ell}},\quad\gamma_{2}=\overline{q}(1+4(1-2^{-1/d})^{-1})^{2\ell}.$
Thus since $(4d(5-2^{-1/d}))^{\frac{2\ell}{1+2\ell}}\leq 20d$ we have the desired result. ∎
4.2 Minimizing the surrogate
Let us now discuss the practical minimization of the surrogate $\mathcal{L}_{1}(g)$ defined in (4.1). We can first observe that $\mathcal{L}_{1}^{1/2}$ is actually a semi-norm, as stated in Proposition 4.11, and thus $\mathcal{L}_{1}$ is a convex function. This is to compare with the functional $\mathcal{J}$ from (2.1) based on Poincaré inequalities, which is highly non-convex in general.
Proposition 4.11.
For any $g\in\mathcal{G}_{1}$ it holds
$\mathcal{L}_{1}(g)=\langle g,g\rangle_{u,1}=\|g\|_{u,1}^{2},$ (4.13)
where $\|\cdot\|_{u,1}$ is the semi-norm associated to the semi-inner product $\langle\cdot,\cdot\rangle_{u,1}$ defined by
$\langle g,h\rangle_{u,1}:=\mathbb{E}\left[\|\nabla u(\mathbf{X})\|_{2}^{2}\langle\nabla g(\mathbf{X}),\Pi^{\perp}_{\nabla u(\mathbf{X})}\nabla h(\mathbf{X})\rangle_{2}\right],\quad\forall g,h\in\mathcal{G}_{1}.$ (4.14)
Now assume that $\mathcal{G}_{1}$ lies in a finite dimensional vector space $\text{span}\{\Phi_{1},\cdots,\Phi_{K}\}$, as in (2.7) for $m=1$. An element $g\in\mathcal{G}_{1}$ admits the representation $g(\mathbf{x})=G^{T}\Phi(\mathbf{x})$ with some $G\in\mathbb{R}^{K}$. Then $\mathcal{L}_{1}(G^{T}\Phi)$ has a simple expression in terms of $G\in\mathbb{R}^{K}$. Indeed, it can be expressed as a positive semi-definite quadratic form of $G$ with some positive semi-definite matrix $H$ which depends on $u$ and $\Phi$, as stated in Proposition 4.12 below.
Proposition 4.12.
For any $G\in\mathbb{R}^{K}$ it holds
$\mathcal{L}_{1}(G^{T}\Phi)=G^{T}HG,$ (4.15)
where $H:=H^{(1)}-H^{(2)}\in\mathbb{R}^{K\times K}$ is a positive semi-definite matrix, with
$H^{(1)}$ $:=\mathbb{E}\left[\|\nabla u(\mathbf{X})\|_{2}^{2}\nabla\Phi(\mathbf{X})^{T}\nabla\Phi(\mathbf{X})\right],$ (4.16)
$H^{(2)}$ $:=\mathbb{E}\left[\nabla\Phi(\mathbf{X})^{T}\nabla u(\mathbf{X})\nabla u(\mathbf{X})^{T}\nabla\Phi(\mathbf{X})\right].$
Proof.
See Section B.4. ∎
Now in order to possibly fulfill the assumptions required to control the suboptimality of $\mathcal{L}_{1}$, we need to restrict the minimization of $\mathcal{L}_{1}$ to a compact subset of $\text{span}\{\Phi_{i}\}_{1\leq i\leq K}$. A convenient such subset is the unit sphere described in (2.8), which we recall in the case $m=1$ is
$\mathcal{G}_{1}=\Big{\{}g:\mathbf{x}\mapsto G^{T}\Phi(\mathbf{x}):G\in\mathbb{R}^{K},G^{T}RG=1\Big{\}},$
with $R\in\mathbb{R}^{K\times K}$ a symmetric positive definite matrix. Recall that the orthogonality condition in (2.8) has no impact on the minimization of $\mathcal{J}$. In addition to that, minimizing $\mathcal{L}_{1}$ over $\mathcal{G}_{1}$ is equivalent to finding the minimal generalized eigenpair of the pair $(H,R)$, as we state in Proposition 4.13.
Proposition 4.13.
The minimizers of $\mathcal{L}_{1}$ over $\mathcal{G}_{1}$ defined in (2.8) are the functions of the form $g^{*}(\mathbf{x})=G^{*T}\Phi(\mathbf{x})$, where $G^{*}$ is a solution to the generalized eigenvalue problem
$\min_{\begin{subarray}{c}G\in\mathbb{R}^{K}\\ G^{T}RG=1\end{subarray}}G^{T}HG,$ (4.17)
with $H$ defined in (4.16).
Let us end this section by discussing on the choice of the matrix $R$. In [3] the authors imposed the condition $\mathbb{E}\left[g(\mathbf{X})^{2}\right]=1$ in order to stabilize the algorithm for minimizing $\mathcal{J}$. This is actually equivalent to taking $R=\mathbb{E}\left[\Phi(\mathbf{X})\Phi(\mathbf{X})^{T}\right]$. However, with our approach, the suboptimality of $\mathcal{L}_{1}$ relies on the control of $\|\nabla g(\mathbf{X})\|^{2}_{2}$ and not $|g(\mathbf{X})|^{2}$. Hence it is more natural to consider a condition of the type $\mathbb{E}\left[\|\nabla g(\mathbf{X})\|_{2}^{2}\right]=1$, which is equivalent to taking $R=\mathbb{E}\left[\nabla\Phi(\mathbf{X})^{T}\nabla\Phi(\mathbf{X})\right]$.
Remark 4.1.
In [22], the authors proposed to select the coefficient matrix $G\in\mathbb{R}^{K\times m}$ whose columns are the eigenvector associated to the largest eigenvalues of
$\tilde{H}=\mathbb{E}\left[\nabla\Phi(\mathbf{X})^{+}\nabla u(\mathbf{X})\nabla u(\mathbf{X})^{T}\nabla\Phi(\mathbf{X})^{+\,T}\right]\in\mathbb{R}^{K\times K},$
where $A^{+}$ denotes the Moore-Penrose pseudo inverse of a matrix $A$ of any size. It turns out that it can be seen as a special case of our approach. First, note that in Definition 2.1, Proposition 2.1 and (4.1) we decided to use the ambient euclidean metric of $\mathbb{R}^{d}$ as Riemannian metric on $\mathcal{X}\subset\mathbb{R}^{d}$. However, we could also formulate it equivalently with any other metric $M:\mathcal{X}\rightarrow\mathbb{R}^{d\times d}$, leading to corresponding loss function $\mathcal{J}^{M}$, convex surrogate $\mathcal{L}_{1}^{M}$ and Poincaré constant $C^{M}(\mathbf{X}|\mathcal{G}_{1})$. Note that this is somehow what is done in [25] where the authors leverage what they call Poincaré inequalities in feature space. Those quantities now involve $\|\cdot\|_{M(\mathbf{X})}^{2}=\langle\cdot,M(\mathbf{X})\cdot\rangle_{2}$ instead of $\|\cdot\|_{2}^{2}$ as well as the orthogonal projection matrices $\Pi^{M(\mathbf{X})}_{\nabla g(\mathbf{X})}$ and $\Pi^{M(\mathbf{X})}_{\nabla u(\mathbf{X})}$ with respect to the inner product matrix $M(\mathbf{X})$. In particular, choosing $M(\mathbf{x})=(\nabla\Phi(\mathbf{x})\nabla\Phi(\mathbf{x})^{T})^{-1}$ yields
$\mathcal{L}_{1}^{M}(G^{T}\Phi)=G^{T}\big{(}\mathbb{E}\left[\|\nabla\Phi(\mathbf{X})^{+}\nabla u(\mathbf{X})\|_{2}^{2}\nabla\Phi(\mathbf{X})^{T}\nabla\Phi(\mathbf{X})^{+\,T}\right]-\tilde{H}\big{)}G,$
Now if we take $R=\mathbb{E}\left[\|\nabla\Phi(\mathbf{X})^{+}\nabla u(\mathbf{X})\|_{2}^{2}\nabla\Phi(\mathbf{X})^{T}\nabla\Phi(\mathbf{X})^{+\,T}\right]$ then we have that
$\min_{\begin{subarray}{c}G\in\mathbb{R}^{K}\\ G^{T}RG=1\end{subarray}}\mathcal{L}^{M}_{1}(G^{T}\Phi)=\min_{\begin{subarray}{c}G\in\mathbb{R}^{K}\\ G^{T}RG=1\end{subarray}}(1-G^{T}\tilde{H}G)=1-\max_{\begin{subarray}{c}G\in\mathbb{R}^{K}\\ G^{T}RG=1\end{subarray}}G^{T}\tilde{H}G.$
As a result, the minimizer of $\mathcal{L}^{M}_{1}$ over $\mathcal{G}_{1}$ is obtained by finding the eigenvector associated to the largest generalized eigenvalue of $(\tilde{H},R)$. Finally, if by any chance $\Phi$ and $u$ are such that $R=I_{K}$, then this eigenvector is also solution to (4.17). In this very specific case, our approach is then equivalent to the one from [22].
5 The case of multiple features
In this section, we propose an extension of the surrogate we introduced in (4.1) to the general case $m\geq 1$. The main problem when trying to do so is that for $m>1$, $\text{span}\{\nabla\}u(\mathbf{X})$ and $\text{span}\{\nabla\}g(\mathbf{X})$ are no longer of the same dimension, thus we can no longer expect to relate $\|\Pi^{\perp}_{\nabla g(\mathbf{X})}\nabla u(\mathbf{x})\|_{2}$ with $\|\Pi^{\perp}_{\nabla u(\mathbf{x})}\nabla g(\mathbf{X})\|_{F}$ in a similar way as in Lemma 4.2. For this reason, the extension we propose relies on learning one feature at a time, in a coordinate descent or greedy approach.
To do so, let us first introduce a splitting of coordinates of the feature map $g\in\mathcal{G}_{m}$. For $1\leq j\leq m$ and $g\in\mathcal{G}_{m}$ we denote by $g_{j}$ and $g_{-j}$ respectively the $j$-th and the other $m-1$ components of $g$, and we define the corresponding sets of functions
$\mathcal{G}_{m}^{(j)}:=\{\mathbf{x}\mapsto h_{j}(\mathbf{x})\,:\,h\in\mathcal{G}_{m}\},\quad\text{and}\quad\mathcal{G}_{m}^{(-j)}:=\{\mathbf{x}\mapsto(h_{i}(\mathbf{x}))_{i\neq j}\,:\,h\in\mathcal{G}_{m}\}.$
Note that, by the invariance property (2.5), $\mathcal{J}(g)$ is invariant to permutations of the coordinates of $g$, thus $\mathcal{J}(g)=\mathcal{J}((g_{-j},g_{j}))$ and minimizing $\mathcal{J}$ on $\mathcal{G}_{m}$ or $\mathcal{G}_{m}^{(j)}\times\mathcal{G}_{m}^{(-j)}$ is equivalent. Note that elements in $\mathcal{G}_{m}^{(j)}$ and $\mathcal{G}_{m}^{(-j)}$ remain coupled via possible constrains imposed in $\mathcal{G}_{m}$. From now on, we consider a fixed $g_{-j}\in\mathcal{G}_{m}^{(-j)}$ and we focus on learning $g_{j}$, meaning minimizing $h\mapsto\mathcal{J}((g_{-j},h))$ on $\mathcal{G}_{m}^{(j)}$. The first thing to notice is that for any matrix $W\in\mathbb{R}^{d\times(m-1)}$ and vectors $\mathbf{w},\mathbf{v}\in\mathbb{R}^{d}$, we can write
$\Pi^{\perp}_{(W|\mathbf{w})}\mathbf{v}=\Pi^{\perp}_{(W|\Pi^{\perp}_{W}\mathbf{w})}\mathbf{v}=\Pi^{\perp}_{\Pi^{\perp}_{W}\mathbf{w}}\Pi^{\perp}_{W}\mathbf{v},$
where $(W|\mathbf{w})\in\mathbb{R}^{d\times m}$ denotes the concatenation of $W$ and $\mathbf{w}$. Hence, by considering
$w_{g,j}(\mathbf{x}):=\Pi^{\perp}_{\nabla g_{-j}(\mathbf{x})}\nabla g_{j}(\mathbf{x})\in\mathbb{R}^{d}\quad\text{and}\quad v_{g,j}(\mathbf{x}):=\Pi^{\perp}_{\nabla g_{-j}(\mathbf{x})}\nabla u(\mathbf{x})\in\mathbb{R}^{d},$ (5.1)
we can write
$\mathcal{J}(g)=\mathbb{E}\left[\|\Pi^{\perp}_{w_{g,j}(\mathbf{X})}v_{g,j}(\mathbf{X})\|_{2}^{2}\right].$
We now end up with an expression for $\mathcal{J}(g)$ which is very similar to what we had in the case $m=1$, where $\nabla g$ and $\nabla u$ are replaced respectively by $w_{g,j}$ and $v_{g,j}$. For this reason, we introduce the following surrogate,
$\mathcal{L}_{m,j}(g):=\mathbb{E}\left[\|v_{g,j}(\mathbf{X})\|_{2}^{2}\|\Pi^{\perp}_{v_{g,j}(\mathbf{X})}w_{g,j}(\mathbf{X})\|_{2}^{2}\right],$ (5.2)
with $w_{g,j},v_{g,j}$ as defined in (5.1). Doing so allows us to make a link between $\mathcal{J}$ and $\mathcal{L}_{m,j}$ by taking $w=w_{g,j}(\mathbf{X})$ and $v=v_{g,j}(\mathbf{X})$ in Lemma 4.2, which yields
$\|w_{g,j}(\mathbf{X})\|_{2}^{2}\|\Pi^{\perp}_{w_{g,j}(\mathbf{X})}v_{g,j}(\mathbf{X})\|_{2}^{2}=\|v_{g,j}(\mathbf{X})\|_{2}^{2}\|\Pi^{\perp}_{v_{g,j}(\mathbf{X})}w_{g,j}(\mathbf{X})\|_{2}^{2}.$ (5.3)
Here again, in the right-hand side we recognize the term within the expectation in $\mathcal{L}_{m,j}$, and in the left-hand side we recognize the term within the expectation in $\mathcal{J}$, up to the factor $\|w_{g,j}(\mathbf{X})\|_{2}^{2}$. A first condition for this to be exploitable is that $\|w_{g,j}(\mathbf{X})\|_{2}^{2}>0$ almost surely. Moreover, since we might want to make use of all the surrogates $\mathcal{L}_{m,1},\cdots,\mathcal{L}_{m,m}$, we will need this condition for all $1\leq j\leq m$. This is actually equivalent to asking for the Assumption 4.3. Finally, we will also rely on a compactness Assumption 5.1 on $\mathcal{G}_{m}$, which is similar to Assumption 4.1 but formulated in the framework of multiple features.
Assumption 5.1.
$\mathcal{G}_{m}$ is a compact subset of $\mathcal{C}^{1}(\mathbb{R}^{d},\mathbb{R}^{m})\cap W^{1,p_{1}}_{\mu}(\mathbb{R}^{d},\mathbb{R}^{m})$ for some $p_{1}>2$, endowed with the norm
$\|g\|=(\|\|g(\mathbf{X})\|^{2}_{2}\|_{p_{1}}^{p_{1}}+\|\|\nabla g(\mathbf{X})\|^{2}_{F}\|_{p_{1}}^{p_{1}})^{1/p_{1}}.$
The structure of this section is as follows. First in Section 5.1 we discuss on the suboptimality of $\mathcal{L}_{m,j}$. Then in Section 5.2 we discuss on the problem of minimizing $\mathcal{L}_{m,j}$.
5.1 suboptimality
Similarly to Proposition 4.5, we can show that in the case $u=f\circ g$ for some $g\in\mathcal{G}_{m}$, minimizing $\mathcal{L}_{m,j}$ is equivalent to minimizing $\mathcal{J}$, as stated in Proposition 5.2 below.
Proposition 5.2.
Assume that $\mathcal{G}_{m}$ satisfies Assumption 4.3, then for all $g\in\mathcal{G}_{m}$ we have $\mathcal{J}(g)=0$ if and only if $\mathcal{L}_{m,j}(g)=0$.
Proof.
See Section C.1
Now again, in general we cannot expect the minimum of $\mathcal{J}$ to be zero, and the ideal setting to quantify suboptimality of $\mathcal{L}_{m,j}$ would be the bi-Lipschitz setting as in Proposition 4.6, which we restate in the context $m\geq 1$ in Proposition 5.3. Note again that such context occurs when considering $\mathcal{G}_{m}$ as in (2.9) with $\mathcal{D}$ containing only bi-Lipschitz diffeomorphisms whose associated constants are uniformly bounded over $\mathcal{D}$. Note also that our analysis in Proposition 5.3 is rather coarse, and with finer assumptions on bi-Lipschitz constant depending on the $j$-th feature, one could probably obtain a sharper bound.
Proposition 5.3.
Assume that there exist $c>0$ and $C<+\infty$ such that for all $g\in\mathcal{G}_{m}$ and $\mathbf{x}\in\mathcal{X}$ it holds $c\leq\sigma_{min}(\nabla g(\mathbf{x}))^{2}\leq\sigma_{max}(\nabla g(\mathbf{x}))^{2}\leq C$. Then we have
$c\mathcal{J}(g)\leq\mathcal{L}_{m,j}(g)\leq C\mathcal{J}(g).$ (5.4)
Thus for fixed $g_{-j}\in\mathcal{G}_{m}^{(-j)}$, if $g_{j}^{*}$ is a minimizer of $h\mapsto\mathcal{L}_{m,j}((g_{-j},h))$ over $\mathcal{G}_{m}^{(j)}$, it yields
$\mathcal{J}((g_{-j},g_{j}^{*}))\leq(C/c)\inf_{h\in\mathcal{G}_{m}^{(j)}}\mathcal{J}((g_{-j},h))$ (5.5)
Proof.
See Section C.2
Without the bi-Lipschitz assumption, the suboptimality of our surrogate $\mathcal{L}_{m}$ should rely on controlling the probability of small and large deviations of $\|w_{g,j}(\mathbf{X})\|_{2}^{2}$, as in the setting $m=1$. For the large deviations, it is rather straightforward since we have $\|w_{g,j}(\mathbf{X})\|_{2}^{2}\leq\|\nabla g_{j}(\mathbf{X})\|_{2}^{2}$, thus we can re-use the results from Section 3. However for the small deviations, meaning controlling the probability of the event $(\|w_{g,j}(\mathbf{X})\|^{2}_{2}\leq\alpha)$ for some $\alpha>0$, the task is much more complicated. The main issue is that the family $\{\nabla g_{i}(\mathbf{X})\}_{1\leq i\leq m}$ is not orthogonal in general.
Still, we can obtain small deviation inequalities, and thus establish a suboptimality result, for the specific case of multiple polynomial features with bounded total degree. If this total degree is $\ell+1$, then we can show that $\mathbf{x}\mapsto\|w_{g,j}(\mathbf{x})\|_{2}^{2}$ is a multivariate rational fraction, whose numerator and denominator have total degree respectively bounded by $2\ell m$ and $2\ell(m-1)$. With this in mind, we can obtain Proposition 5.4.
Proposition 5.4.
Assume that $\mathbf{X}$ is an absolutely continuous random variable on $\mathbb{R}^{d}$ whose distribution is $s$-concave with $s\in(0,1/d]$. Assume $\mathcal{G}_{m}$ is a compact set of polynomials with total degree at most $\ell+1$ and which satisfies Assumption 4.3. Then for all $g\in\mathcal{G}_{m}$ and $1\leq j\leq m$ it holds
$\mathbb{P}\left[\|\Pi^{\perp}_{\nabla g_{-j}}\nabla g_{j}(\mathbf{X})\|^{2}_{2}\leq\underline{q}\varepsilon\right]\leq\eta\varepsilon^{1/2\ell m},$ (5.6)
with $\underline{q}>0$ and $\eta=\eta_{s,\mathcal{G}_{m}}^{small}<+\infty$ both defined in (C.3).
Proof.
See Section C.3. ∎
Note that the quantities $\underline{q}$ and $\eta$ involved in Proposition 5.4 are not the same as in Proposition 3.3. In particular, although $\underline{q}$ is still a uniform lower bound on some set of quantiles, the quantiles involved are not the ones of $\|w_{g,j}(\mathbf{X})\|$. Note also that the rate of decay of the small deviation probability is much worse than in the one feature case. Indeed, here, this decay scales with both the polynomial degree $\ell$ and the number of features $m$.
Finally, let us end this section by stating in Proposition 5.5 below a suboptimality result on the minimizer of our surrogate $h\mapsto\mathcal{L}_{m,j}((g_{-j},h))$, which relies on large and small deviation inequalities on $\|\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}\nabla g_{j}(\mathbf{X})\|_{2}^{2}$. Note that there are two main differences between the suboptimality result for $m=1$ feature in Corollary 4.10 and the suboptimality result for $m>1$ features in Proposition 5.5. The first difference is that for multiple features, the result actually concerns only one of the $m$ features. The second is that the exponent is now $1/(1+2\ell m)$ instead of $1/(1+2\ell)$, which is much worse. Additionally, the larger is $m$, in other words the more features we have to learn, the worse is the suboptimality result.
Proposition 5.5.
Assume that $\mathbf{X}$ is an absolutely continuous random variable on $\mathbb{R}^{d}$ whose distribution is $s$-concave with $s\in(0,1/d]$. Assume $\mathcal{G}_{m}$ is a compact set of polynomial with total degree at most $\ell+1$ and which satisfies Assumption 4.3. Then, for fixed $g_{-j}\in\mathcal{G}_{m}^{(-j)}$, if $g_{j}^{*}$ is a minimizer of $h\mapsto\mathcal{L}_{m,j}((g_{-j},h))$ over $\mathcal{G}_{m}^{(j)}$, it holds
$\mathcal{J}((g_{-j},g_{j}^{*}))\leq\tilde{\gamma}_{3}\inf_{h\in\mathcal{G}_{m}^{(j)}}\mathcal{J}((g_{-j},h))^{\frac{1}{1+2\ell m}},$ (5.7)
with $\tilde{\gamma}_{3}<+\infty$ defined at the end of Section C.4.
Proof.
See Section C.4. ∎
5.2 Minimizing the surrogate
Now we discuss on the practical minimization of the surrogate $\mathcal{L}_{m,j}$ introduced in (5.2). At first glance, $\mathcal{L}_{m,j}$ does not have much better properties than $\mathcal{J}$ since it involves $\Pi_{\nabla g_{-j}(\mathbf{X})}$. However, optimizing the $j$-th coordinate is actually similar to the one feature case. Indeed for a fixed $g_{-j}\in\mathcal{G}_{-j}$ the function
$\mathcal{G}_{m}^{(j)}\ni h\mapsto\mathcal{L}_{m,j}((g_{-j},h))$
has the same properties as $\mathcal{L}_{1}$, in other words it is a squared semi-norm, as stated in Proposition 5.6. This is to compare with the function $h\mapsto\mathcal{J}((g_{-j},h))$ which is highly non-convex in general. It is important to note that the associated semi-norm depends on both $u$ and $g_{-j}$.
Proposition 5.6.
For any $g\in\mathcal{G}_{m}$ it holds
$\mathcal{L}_{m,j}(g)=\langle g_{j},g_{j}\rangle_{u,m,j}=\|g_{j}\|_{u,m,j}^{2},$ (5.8)
where $\|\cdot\|_{u,m,j}$ is the semi-norm associated to the semi-inner product $\langle\cdot,\cdot\rangle_{u,m,j}$ defined by
$\langle h,l\rangle_{u,m,j}:=\mathbb{E}\left[\|v_{g,j}(\mathbf{X})\|_{2}^{2}\langle\nabla h(\mathbf{X}),\Pi^{\perp}_{v_{g,j}(\mathbf{X})}\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}\nabla l(\mathbf{X})\rangle_{2}\right],\quad\forall h,l\in\mathcal{G}_{m}^{(j)},$ (5.9)
with $w_{g,j}$ and $v_{g,j}$ as in (5.1).
Proof.
The projectors $\Pi^{\perp}_{v_{g,j}(\mathbf{X})}$ and $\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}$ commute since $v_{g,j}(\mathbf{X})\in\text{span}\{(\}\nabla g_{-j}(\mathbf{X}))^{\perp}$. The definition of $\mathcal{L}_{m,j}$ in (5.2) then yields the result. ∎
Now again when $\mathcal{G}_{m}^{(j)}\subset\text{span}\{\Phi_{i}\}_{1\leq i\leq K}$ we can express $G_{j}\mapsto\mathcal{L}_{m,j}((g_{-j},G_{j}^{T}\Phi))$ as a positive semi-definite quadratic form on $G_{j}$ with some positive semi-definite matrix $H_{m,j}$ which depends on $u$, $\Phi$ and $g_{-j}$, as stated in Proposition 5.7 below.
Proposition 5.7.
Let $g_{-j}\in\mathcal{G}_{m}^{(-j)}$ be fixed. For any $G_{j}\in\mathbb{R}^{K}$ and $g_{j}(\mathbf{x})=G_{j}^{T}\Phi(\mathbf{x})$, it holds
$\mathcal{L}_{m,j}((g_{-j},g_{j}))=G_{j}^{T}H_{m,j}G_{j},$ (5.10)
where $H_{m,j}=H_{m,j}^{(1)}-H_{m,j}^{(2)}\in\mathbb{R}^{K\times K}$ is a positive semi-definite matrix with
$H_{m,j}^{(1)}$ $:=\mathbb{E}\left[\|\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}\nabla u(\mathbf{X})\|_{2}^{2}\nabla\Phi(\mathbf{X})^{T}\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}\nabla\Phi(\mathbf{X})\right],$ (5.11)
$H_{m,j}^{(2)}$ $:=\mathbb{E}\left[\|\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}\nabla u(\mathbf{X})\|_{2}^{2}\nabla\Phi(\mathbf{X})^{T}\nabla u(\mathbf{X})\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}\nabla u(\mathbf{X})^{T}\nabla\Phi(\mathbf{X})\right].$
Proof.
See Section C.5. ∎
Now, let us study the case when $\mathcal{G}_{m}$ has the form (2.8). In this case, we have a constrained minimization of a Rayleigh quotient, where the constrains are $G_{i}^{T}RG_{j}=\delta_{i,j}$. It turns out that this minimization problem is equivalent to some generalized eigenvalue problem, with orthogonality matrix $R$. Indeed, be noticing that $H_{m,j}G_{-j}=0$, we have for any $\alpha$ that $H_{m,j}+\alpha RG_{-j}G_{-j}^{T}R$ admits $G_{-j}$ as generalized eigenvector with associated eigenvalue $\alpha$. Now to be sure that this is not the smallest one, we shall take $\alpha$ large enough, in particular larger than any generalized eigenvalue $\lambda$ of $(H_{m,j},R)$. Note that by definition of $\mathcal{L}_{m,j}$ in (5.2), it holds
$G_{j}H_{m,j}G_{j}\leq\mathbb{E}\left[\|\nabla u(\mathbf{X})\|_{2}^{2}\|\nabla g_{j}(\mathbf{X})\|_{2}^{2}\right]=G_{j}^{T}H^{(1)}G_{j},$
with $H^{(1)}$ defined in Equation 4.16. Hence choosing $\alpha$ as the largest generalized eigenvalue of $(H^{(1)},R)$ yields that $G_{-j}$ is not associated to the smallest generalized eigenvector of $(H_{m,j}+\alpha RG_{-j}G_{-j}^{T}R,R)$, and that the corresponding generalized eigenvector is orthogonal to $G_{-j}$ with respect to $R$. This is summarized in Proposition 5.8 below.
Proposition 5.8.
Let $g_{-j}=G_{-j}^{T}\Phi\in\mathcal{G}_{m}^{(-j)}$. The minimizers of $h\mapsto\mathcal{L}_{m,j}((g_{-j},h))$ over $\mathcal{G}_{m}^{(j)}$ defined in (2.8) are the functions of the form $g_{j}^{*}=G_{j}^{*\,T}\Phi$, where $G_{j}^{*}$ is a solution to the generalized eigenvalue problem
$\min_{\begin{subarray}{c}v\in\mathbb{R}^{K}\\ v^{T}Rv=1\end{subarray}}v^{T}(H_{m,j}+\alpha H_{m,j}^{(3)})v$ (5.12)
with $H_{m,j}$ as defined in (5.11), $H_{m,j}^{(3)}:=RG_{-j}G_{-j}^{T}R$ and $\alpha$ as prescribed above.
With Proposition 5.8, we have an explicit formulation to compute the $j$-th feature $g_{j}=G_{j}^{T}\Phi$ from the $j-1$ previously learnt features $g_{-j}=G_{-j}^{T}\Phi$, based on the surrogate we introduced in (5.2). We now propose to leverage it in a greedy algorithm, in which we compute successively $G_{1},\cdots,G_{m}$. The algorithm is described in Algorithm 1. Note that in practice, at the end of each greedy iteration, one may consider running a few iterations of some minimization algorithm on $G\mapsto\mathcal{J}(G^{T}\Phi)$. Indeed, one shall keep in mind that the purpose of our surrogates $\mathcal{L}_{1}$ and $\mathcal{L}_{m,j}$ is to obtain a suboptimal solution to the minimization of $\mathcal{J}$. Although the latter optimization problem is rather difficult to solve, we can still leverage first and second order optimization algorithms, as proposed in [3].
Algorithm 1 Greedy algorithm to learn multiple features
1:Number of features $m\geq 2$, basis $\Phi:\mathbb{R}^{d}\mapsto\mathbb{R}^{K}$ of feature space, inner product matrix $R\in\mathbb{R}^{K\times K}$.
2:$G\in\mathbb{R}^{k\times m}$ and $G^{T}RG=I_{m}$.
3:Compute $H^{(1)},H^{(2)}\in\mathbb{R}^{K\times K}$ defined in (4.16).
4:Compute $G_{1}\in\mathbb{R}^{K}$ solution to (4.17) and $\alpha$ largest generalized eigenvalue of $(H^{(1)},R)$.
5:Set $G\leftarrow G_{1}$.
6:for $2\leq j\leq m$ do
7:     Compute $H_{j,j}^{(1)},H_{j,j}^{(2)},H_{j,j}^{(3)}\in\mathbb{R}^{K\times K}$ from (5.11) and (5.12) with $g_{-j}=G^{T}\Phi$.
8:     Compute $G_{j}$ solution to (5.12).
9:     Set $G\leftarrow(G_{1}|\cdots|G_{j-1}|G_{j})\in\mathbb{R}^{K\times j}$.
10:end for
6 Numerical experiments
6.1 Setting
Benchmark functions.
We consider the following four benchmark functions $u_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ with $d=8$:
$u_{1}(\mathbf{x}):=\sin(\frac{4}{\pi^{2}}\mathbf{x}^{T}\mathbf{x}),$ $\quad u_{2}(\mathbf{x}):=\cos(\frac{1}{2}\mathbf{x}^{T}\mathbf{x})+\sin(\frac{1}{2}\mathbf{x}^{T}M\mathbf{x}),$
$u_{3}(\mathbf{x}):=\exp(\frac{1}{d}\sum_{i=1}^{d}\sin(x_{i})e^{\cos(x_{i})}),$ $\quad u_{4}(\mathbf{x}):=\frac{2\pi x_{3}(x_{4}-x_{6})}{\ln(x_{2}/x_{1})(1+\frac{2x_{7}x_{3}}{\ln(x_{2}/x_{1})x_{1}^{2}x_{8}}+\frac{x_{3}}{x_{5}}).}$
In $u_{2}$, we choose the matrix $M=(\frac{1}{i+j-1})_{1\leq i,j\leq d}\in\mathbb{R}^{d\times d}$. For $u_{1},u_{2}$ and $u_{3}$ we take $\mathbf{X}\sim\mathcal{U}(]-\frac{\pi}{2},\frac{\pi}{2}[^{d})$. For the so called Borehole function $u_{4}$ we take $X_{1}\sim\mathcal{N}(0.1,0.0161812)$, $X_{2}\sim\log\mathcal{N}(7.71,1.0056)$, $X_{3}\sim\mathcal{U}(63070,115600)$, $X_{4}\sim\mathcal{U}(990,1110)$, $X_{5}\sim\mathcal{U}(63.1,116)$, $X_{6}\sim\mathcal{U}(700,820)$, $X_{7}\sim\mathcal{U}(1120,1680)$ and $X_{8}\sim\mathcal{U}(9855,12045)$.
The benchmark $u_{1}$ illustrates Proposition 4.5, as it can be expressed as a univariate function of a polynomial, while not being a polynomial. Thus taking $\mathcal{G}_{1}$ as a suitable set of polynomials, we know that $\inf_{\mathcal{G}_{1}}\mathcal{J}=0$, and in that case the minimizers of $\mathcal{J}$ and $\mathcal{L}_{1}$ are the same. Hence $u_{1}$ is perfectly suited for the surrogate we introduced in Section 4. Similarly, the benchmark $u_{2}$ is such that $\inf_{\mathcal{G}_{2}}\mathcal{J}=0$ for a suitable set of polynomials $\mathcal{G}_{2}$. Hence $u_{2}$ seems well suited for the surrogate we introduced in Section 5, although we have less theoretical guarantees. Finally, the benchmarks $u_{3}$ and $u_{4}$ were investigated in the context of nonlinear dimension reduction respectively in [25] and [3].
Quantities monitored.
In our experiments, we will monitor $4$ quantities. The first two are the Poincaré inequality based quantity $\mathcal{J}(g)$ defined in (2.1) and the final approximation error $e_{g}(f)$ defined by
$e_{g}(f):=\|u-f\circ g\|_{L^{2}}.$
We estimate these quantities with their Monte-Carlo estimators on test samples $\Xi^{test}\subset\mathcal{X}$ of sizes $N^{test}=1000$, not used for learning. We also monitor the Monte-Carlo estimators $\widehat{\mathcal{J}}(g)$ and $\widehat{e}_{g}(f)$ on some training set $\Xi^{train}\subset\mathcal{X}$ of various sizes $N^{train}$, which will be the quantities directly minimized to compute $g$ and $f$. More precisely, we draw $20$ realizations of $\Xi^{train}$ and $\Xi^{test}$ and monitor the quantiles of those $4$ quantities over those $20$ realizations.
Sets of feature maps.
We consider feature maps of the form (2.8) with $\Phi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{K}$ a multivariate polynomial obtained by sparse tensorizations of univariate orthonormal polynomials. Those are obtained by selecting multi-indices with $p\in[0,+\infty]$ norm bounded by $k\in[1,+\infty)$, where we see $(p,k)$ as hyperparameter that we will optimize using a cross-validation procedure.
More precisely, for a product measure $d\mu(\mathbf{x})=\bigotimes_{\nu=1}^{d}d\mu_{\nu}(x_{\nu})$ and $1\leq\nu\leq d$, we consider $(\phi_{i}^{\nu})_{i\in\mathbb{N}}$ an orthonormal polynomial basis of $L^{2}_{\mu_{\nu}}$, meaning $\mathbb{E}\left[\phi_{i}^{\nu}(X_{\nu})\phi_{j}^{\nu}(X_{\nu})\right]=\delta_{i,j}$, and $\phi_{i}^{\nu}$ has degree $i$. Then $(\phi_{\alpha})_{\alpha\in\mathbb{N}^{d}}$ is an orthonormal basis of $L^{2}_{\mu_{\nu}}$, where for a multi-index $\alpha\in\mathbb{N}^{d}$ we define the multivariate polynomial $\phi_{\alpha}$ as
$\phi_{\alpha}(\mathbf{x}):=\prod_{\nu=1}^{d}\phi_{\alpha_{\nu}}^{\nu}(x_{\nu}).$
The coordinates of $\Phi$ are then given by multi-indices from some prescribed set $\Lambda\subset\mathbb{N}^{d}$. Here we consider sets of multi-indices of the form $\Lambda_{p,k}:=\{\alpha\in\mathbb{N}^{d}:\|\alpha\|_{p}\leq k\}\setminus\{0\}$, which yields an associated basis $\Phi=\Phi^{\Lambda_{p,k}}$. Note that we excluded the multi-index 00 so that Assumption 4.3 is satisfied. Note also that such definition ensures that $id\in\text{span}\{\Phi_{1},\cdots,\Phi_{K}\}$, thus $\mathcal{G}_{m}$ contains all linear feature maps, including the one corresponding to the active subspace method. The hyperparameters $(p,k)$ are selected using a cross-validation procedure described bellow.
Computing feature maps.
For fixed hyperparameters $(p,k)$, associated basis $\Phi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{K}$ and a fixed training set $\Xi^{train}\subset\mathcal{X}$ of size $N^{train}$, we aim to minimize the Monte-Carlo estimation $\mathbb{R}^{K\times m}\ni G\mapsto\widehat{\mathcal{J}}(G^{T}\Phi)$ evaluated from the samples in $\Xi^{train}$. To do so, we consider $3$ learning procedures.
The first procedure, which we consider as the reference, is based on a preconditioned nonlinear conjugate gradient algorithm on the Grassmann manifold $Grass(m,K)$. We use $\hat{\Sigma}(G)\in\mathbb{R}^{K\times m}$ as preconditioning matrix at point $G\in\mathbb{R}^{K\times m}$, which is the Monte-Carlo estimation of $\Sigma(G)$ defined in [3, Proposition 3.2]. We choose as initial point the matrix $G^{0}\in\mathbb{R}^{K\times m}$ which minimizes $\widehat{\mathcal{J}}$ on the set of linear features, which corresponds to the active subspace method. We denote this reference procedure as GLI, standing for Grassmann Linear Initialization. The second procedure is based on Algorithm 1, where the surrogates $\mathcal{L}_{m,j}$ are estimated via Monte-Carlo based on $\Xi^{train}$. Within Algorithm 1 we use the orthogonality matrix $R=\mathbb{E}\left[\nabla\Phi(\mathbf{X})^{T}\nabla\Phi(\mathbf{X})\right]$. We denote this procedure as SUR, standing for SURrogate. The third procedure is the same as the first one, but where the initial point is given by the second procedure. We denote this procedure as GSI, standing for Grassmann Surrogate Initialization. Note that after procedure, we orthonormalize again $g$ such that $\mathbb{E}\left[\nabla g(\mathbf{X})^{T}\nabla g(\mathbf{X})\right]=I_{m}$ in order to obtain better stability of the subsequent regression problem.
Finally, for each procedure, we select specific hyperparameters $(p,k)$ using a $5$-fold cross-validation procedure on $\Xi^{train}$, among the $10$ following values,
$(p,k)\in\big{(}\{0.8\}\times\{2,3,4,5\}\big{)}\cup\big{(}\{0.9\}\times\{2,3,4\}\big{)}\cup\big{(}\{1\}\times\{1,2,3\}\big{)}.$
Regression function.
Once we have learnt $g\in\mathcal{G}_{m}$, we then perform a classical regression task to learn a regression function $f:\mathbb{R}^{m}\rightarrow\mathbb{R}$, with $g(\mathbf{X})\in\mathbb{R}^{m}$ as input variable and $u(\mathbf{X})\in\mathbb{R}$ as output variable. In particular here, we have chosen to use kernel ridge regression with gaussian kernel $\kappa(\mathbf{y},\mathbf{z}):=\exp(-\gamma\|\mathbf{y}-\mathbf{z}\|^{2}_{2})$ for any $\mathbf{y},\mathbf{z}\in\mathbb{R}^{m}$ and some hyperparameter $\gamma>0$. Then with $\{\mathbf{z}^{(i)}\}_{1\leq i\leq N^{train}}:=g(\Xi^{train})\subset\mathbb{R}^{m}$, we consider
$f:\mathbf{z}\mapsto\sum_{i=1}^{N^{train}}a_{i}\kappa(\mathbf{z}^{(i)},\mathbf{z}),$
with $\mathbf{a}:=(K+\alpha I_{N})^{-1}\mathbf{u}\in\mathbb{R}^{N^{train}}$ for some regularization parameter $\alpha>0$, where $K:=(\kappa(\mathbf{z}^{(i)},\mathbf{z}^{(j)}))_{1\leq i,j\leq N^{train}}$ and $\mathbf{u}:=(u(\mathbf{x}^{(i)}))_{1\leq i\leq N^{train}}$. Here the kernel parameter $\gamma$ and the regularization parameter $\alpha$ are hyperparameters learnt using a $10$-fold cross-validation procedure.
Note that we followed a procedure similar to [3, Algorithm 4]. In particular the hyperparameters for $g$ and $f$ are fitted via two separate cross-validation procedures. One may also consider optimizing both in one global cross-validation procedure, although it would be much more costly.
Implementation.
The cross-validation procedures as well as the Kernel ridge regression rely on the library sklearn [20]. The optimization on Grassmann manifolds rely on on the library pymanopt [24]. The orthonormal polynomials feature maps rely on the python library tensap [19]. Open-source implementation of the approach proposed in this work, as well as tutorials, are available at https://github.com/alexandre-pasco/tensap/tree/poincare_learning.
6.2 The case of a single feature
In this section we investigate the performances of the three learning procedures GLI, SUR and GLS when learning $m=1$ feature to approximate $u_{1}$, $u_{3}$ and $u_{4}$.
Figure 1: Evolution of quantiles with respect to the size of the training sample for $u_{1}$ with $m=1$. The quantiles $50\%$, $90\%$ and $100\%$ are represented respectively by the continuous, dashed and dotted lines.
Let us start with $u_{1}$ for which results are displayed in Figure 1. Recall that this benchmark function can be exactly expressed as a function of $m=1$ polynomial feature of total degree $2$.
Firstly for all $N^{train}$, we observe that SUR and GSI always yield the minimizer of $\widehat{\mathcal{J}}$, which turns out to be 00 as for $\mathcal{J}$. This is the reason why they are not displayed in the first two plots, as they are equal to machine zero, and why SUR and GSI yield the same results, as the minimum is already achieved with SUR. This could be expected from Proposition 4.5, although it would probably have failed for even smaller $N^{train}$. As a result, we know that the minimum of $e_{g}$ over all possible regression functions is 00. However, we still observe that the approximation errors $e_{g}(f)$ may remain rather high, in particular for $u_{2}$. This translates the fact that the model class we choose for $f$ is not suitable to approximate $\mathbf{z}\mapsto\mathbb{E}\left[u(\mathbf{X})|g(\mathbf{X})=\mathbf{z}\right]$, especially for small training sets. In particular, since $\widehat{e}_{g}(f)$ is rather small, it seems that $g(\Xi^{train})$ does not contain enough points located near the borders of the set $g(\mathcal{X})$. This is most probably due to the fact that the sampling strategy used to sample $\mathbf{X}$ is not a good sampling strategy to sample $g(\mathbf{X})$, and maybe a re-sampling with a more suitable strategy would help circumventing the problem.
Secondly for small $N^{train}$, we observe that GLI sometimes fails to converge towards the minimum of $\widehat{\mathcal{J}}$. When it happens, we may not learn good functions $f$, that leads to large approximation errors $\widehat{e}_{g}(f)$ and $e_{g}(f)$.
Figure 2: Evolution of quantiles with respect to the size of the training sample for $u_{3}$ with $m=1$. The quantiles $50\%$, $90\%$ and $100\%$ are represented respectively by the continuous, dashed and dotted lines.
Let us continue with $u_{3}$, whose associated results are displayed in Figure 2. Firstly for all $N^{train}$, we observe that SUR and GSI perform at least as good as GLI, regarding all 4 monitored quantities. Also, SUR and GSI yield very similar results, which can be interpreted as the minimizer of ^$\widehat{\mathcal{L}}_{1}$ being near a local minimum of $\widehat{\mathcal{J}}$, if not near a global one.
Secondly for rather small $N^{train}$, we observe that SUR and GSI perform better than GLI on all $4$ quantities monitored. We also observe that the results obtained with SUR and GSI are much less varying than with GLI over the $20$ realizations, especially regarding $\widehat{\mathcal{J}}(g)$ and $\mathcal{J}(g)$.
Figure 3: Evolution of quantiles with respect to the size of the training sample for $u_{4}$ with $m=1$. The quantiles $50\%$, $90\%$ and $100\%$ are represented respectively by the continuous, dashed and dotted lines.
Let us finish with $u_{4}$, whose associated results are displayed in Figure 3. Firstly for all $N^{train}$, we again observe that SUR and GSI yield very similar results. We also observe on the $100\%$ quantile that all 3 methods suffer from a rather significantly varying performance over the $20$ realizations, and that no method is always better than the others. For example for $N^{train}=100$, the worst case errors for GLI are significantly better than with SUR and GSI, and vice versa for $N^{train}=75$. However for $N^{train}\geq 75$, we observe on the $50\%$ and $90\%$ quantiles that GLI, SUR and GSI perform similarly.
Secondly for $N^{train}=50$, we observe that $\widehat{\mathcal{J}}(g)$ is often much lower for SUR and GSI than for GLI, while there is not much of a gap on $\mathcal{J}(g)$. This means that, although SUR and GSI performed much better than GLI when it comes to minimizing $\widehat{\mathcal{J}}$, the minimizer of the latter is rather far from the one of $\mathcal{J}$. Such behavior is often referred to as overfitting.
6.3 The case of multiple features
In this section we investigate the performances of the three learning procedures GLI, SUR and GLS when learning $m=2$ features to approximate $u_{2}$, $u_{3}$ and $u_{4}$.
Figure 4: Evolution of quantiles with respect to the size of the training sample for $u_{2}$ with $m=2$. The quantiles $50\%$, $90\%$ and $100\%$ are represented respectively by the continuous, dashed and dotted lines.
Let us start with $u_{2}$ for which results are displayed in Figure 4. Recall that this benchmark can be exactly expressed as a bivariate function of $m=2$ polynomial features of total degree $2$.
Firstly for all $N^{train}$, we observe that GSI outperforms GLI in all aspects, although the gap shrinks for large $N^{train}$. In particular, it often yields the minimizer of $\widehat{\mathcal{J}}$, which turns out to be 00 as for $\mathcal{J}$. We also observe that SUR and GSI now yield rather different results, which could be expected since the suboptimality results of our surrogates are much weaker for $m>1$ than for $m=1$. In particular, SUR performs poorly here, although it appears that it yields a good initialization point for GSI.
Secondly for small $N^{train}$, we observe that GLI often fails to yield the minimizer of $\widehat{\mathcal{J}}$ while GSI is often successful, although there were realizations where GSI also failed.
Thirdly for large $N^{train}$, we observe that GLI still sometimes fails to yield the minimizer of $\widehat{\mathcal{J}}$ while GSI is always successful.
Figure 5: Evolution of quantiles with respect to the size of the training sample for $u_{3}$ with $m=2$. The quantiles $50\%$, $90\%$ and $100\%$ are represented respectively by the continuous, dashed and dotted lines.
Figure 6: Evolution of quantiles with respect to the size of the training sample for $u_{4}$ with $m=2$. The quantiles $50\%$, $90\%$ and $100\%$ are represented respectively by the continuous, dashed and dotted lines.
Let us continue with $u_{3}$ and $u_{4}$, whose associated results are displayed respectively in Figures 5 and 5.
Firstly we observe when comparing Figures 2 and 6 that adding a feature did not improve much, if not deteriorated, the performances regarding $\mathcal{J}(g)$, $\widehat{e}_{g}(f)$ and $e_{g}(f)$, for all three procedures.
Secondly we observe that for rather small values of $N^{train}$, GSI performs much better at minimizing $\widehat{\mathcal{J}}(g)$ than SUR and GLI. In particular for $u_{3}$ with $N^{train}=75$ and for $u_{4}$ with $N^{train}=50$, GSI mostly yields machine zero values for $\widehat{\mathcal{J}}(g)$. However, the corresponding values for $\mathcal{J}(g)$ are much larger, which again translates same overfitting behavior. This results in rather large approximation errors $\widehat{e}_{g}(f)$ and $e_{g}(f)$. Hence for those benchmarks with $m=2$, it appears that the approaches SUR and GSI we proposed did not provide clear performance gain.
6.4 Conclusion
From what we observed in Sections 6.2 and 6.3, we can draw $3$ general observations.
Firstly, GSI mostly performed better than GLI for minimizing $\widehat{\mathcal{J}}(g)$, which is the Monte-Carlo estimation of $\mathcal{J}$ using training samples. This is especially true for rather (but not too) small training sample sizes $N^{train}$ For larger $N^{train}$ however, both GSI and GLI performed similarly. Now the problem is that when $N^{train}$ is too small, the minimizer of $\widehat{\mathcal{J}}$ may be rather far from the one of $\mathcal{J}$, which is usually denoted as an overfitting behavior. In such cases, the performance of GSI is similar to GLI, or even deteriorated. We especially observed this overfitting behavior for $m=2$ on benchmarks $u_{3}$ and $u_{4}$ where the minimum of $\mathcal{J}$ was rather large.
Secondly, for $m=1$ feature, SUR yielded features that seemed to be near local minima of $\widehat{\mathcal{J}}$, and GSI did not yield much smaller values. In particular when the minimum of $\widehat{\mathcal{J}}$ was 00, as for $u_{1}$ and $u_{2}$, SUR always found it, while there were realizations where GLI did not. In general, our observations suggest that our approach SUR, or its variant GSI, mostly learn significantly better features for (not too) small training sets.
Thirdly, for $m=2$ features, SUR yielded features that seemed rather far from local minima of $\widehat{\mathcal{J}}$, while GSI yielded significantly lower values. When the problem was reducible to $m=2$ dimensions, GSI often yielded the correct dimension reduction, contrarily to GLI. However for benchmarks with not enough potential dimensionality reduction, it resulted in a rather strong overfitting behavior, resulting in mitigated results on $\mathcal{J}(f)$ and $e_{g}(f)$.
Appendix A Proofs for Section 3
A.1 Proof of Proposition 3.3
For readability, in this section we drop the subscript $\mathcal{K}$ in the expressions $k_{\mathcal{K}},A_{\mathcal{K}}$ and $\underline{q}_{\mathcal{K}}$. Proposition 3.3 is essentially based on the small deviation inequality from [10], which states that a function $h:\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfying (3.2) with constants $k$ and $A$ satisfies
$\mathbb{P}\left[|h(\mathbf{X})|^{1/k}\leq Q_{|h(\mathbf{X})|^{1/k}}(1/2)\varepsilon\right]\leq A(1-2^{-s})s^{-1}\varepsilon,\quad\forall\varepsilon\in(0,1).$
Using the fact that $Q_{|h(\mathbf{X})|^{1/k}}(1/2)^{k}=Q_{|h(\mathbf{X})|}(1/2)\geq\underline{q}$, we thus have
$\mathbb{P}\left[|h(\mathbf{X})|\leq\underline{q}\varepsilon\right]\leq\mathbb{P}\left[|h(\mathbf{X})|\leq Q_{|h(\mathbf{X})|}(1/2)\varepsilon\right]\leq A(1-2^{-s})s^{-1}\varepsilon^{1/k},\quad\forall\varepsilon\in(0,1).$
Finally, taking $\eta=\eta^{small}_{s,\mathcal{K}}=\max(1,(1-2^{-s})s^{-1}A)$ and observing that $\eta\varepsilon^{1/k}\geq 1$ whenever $\varepsilon\geq 1$ yields the desired result.
A.2 Proof of Proposition 3.4
For readability, in this section we drop the subscript $\mathcal{K}$ in the expressions $k_{\mathcal{K}},A_{\mathcal{K}}$ and $\overline{q}_{\mathcal{K}}$. Proposition 3.4 is essentially based on the large deviation inequality in [10], which states that a function $h:\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfying (3.2) with constants $k$ and $A$ satisfies
$\forall t>1,\quad\mathbb{P}\left[|h(\mathbf{X})|^{1/k}\geq AQ_{|h(\mathbf{X})|^{1/k}}(1/2)t\right]\leq\left\{\begin{aligned} (1-(1-2^{-s})t)^{1/s}_{+},\quad&s\in(0,1/d],\\ \exp(-\log(2)t)\quad&s=0,\\ (2^{-s}-1)^{1/s}t^{1/s}\quad&s<0.\end{aligned}\right.$
Using $Q_{|h(\mathbf{X})|^{1/k}}(1/2)^{k}=Q_{|h(\mathbf{X})|}(1/2)\leq\overline{q}$ and substituting $t$ by $A^{-1}t^{1/k}$ yields
$\forall t>A^{k},\quad\mathbb{P}\left[|h(\mathbf{X})|\geq\overline{q}t\right]\leq\left\{\begin{aligned} (1-(1-2^{-s})A^{-1}t^{1/k})^{1/s}_{+},\quad&s\in(0,1/d],\\ \exp(-\log(2)A^{-1}t^{1/k})\quad&s=0,\\ (2^{-s}-1)^{1/s}A^{-1}t^{1/sk}\quad&s<0.\end{aligned}\right.$ (A.1)
Now, we will show that the previous equation actually holds for all $t\geq 1$, then we will weaken the bound so that it holds for all $t>0$. Recall that by definition $A\geq 1$.
Case $0<s\leq 1/d$.
Let $t\in[1,A^{k}]$. Since $x\mapsto\mathbb{P}\left[|h(\mathbf{X})|\geq x\right]$ is decreasing and $t\geq 1$, we have that
$\mathbb{P}\left[|h(\mathbf{X})|\geq\overline{q}t\right]\leq\mathbb{P}\left[|h(\mathbf{X})|\geq\overline{q}\right]\leq\mathbb{P}\left[|h(\mathbf{X})|\geq Q_{|h(\mathbf{X})|}(1/2)\right]=1/2.$
Now since $x\mapsto(1-(1-2^{-s})A^{-1}x^{1/k})^{1/s}_{+}$ is decreasing and $t\leq A^{k}$, we have that
$(1-(1-2^{-s})A^{-1}t^{1/k})^{1/s}_{+}\geq(1-(1-2^{-s}))^{1/s}_{+}=1/2,$
which yields that (A.1) holds for all $t\geq 1$ when $s>0$. Now let $t>0$ and denote $\eta=\eta^{large}_{s,\mathcal{K}}:=A(1-2^{-s})^{-1}$. We have that $(1-t^{1/k}/\eta)\leq(1-(t^{1/k}-1)/\eta)$, and if $t\in(0,1]$ we have that $1\leq(1-(t^{1/k}-1)/\eta)$. Thus, we have the desired result for $0<s\leq 1/d$,
$\forall t>0,\quad\mathbb{P}\left[|h(\mathbf{X})|\geq\overline{q}t\right]\leq(1-(t^{1/k}-1)/\eta).$
Case $s=0$.
Let $t\in[1,A^{k}]$. As for the case $s>0$, we have $\mathbb{P}\left[|h(\mathbf{X})|\geq\overline{q}t\right]\leq 1/2$. Now since $x\mapsto\exp(-\log(2)A^{-t}x^{1/k})$ is decreasing and $t\leq A^{k}$, we have that
$\exp(-\log(2)A^{-t}t^{1/k})\geq\exp(-\log(2))=1/2,$
which yields that (A.1) holds for all $t\geq 1$ when $s=0$. Now, let $t>0$ and denote $\eta=\eta^{large}_{s,\mathcal{K}}:=A/\log(2)$. We have that $\exp(-t^{1/k}/\eta)\leq\exp(-(t^{1/k}-1)/\eta)$ and if $t\in(0,1]$ we have that $1\leq\exp(-(t^{1/k}-1)/\eta)$. Thus, we have the desired result for $s=0$,
$\forall t>0,\quad\mathbb{P}\left[|h(\mathbf{X})|\geq\overline{q}t\right]\leq\exp(-(t^{1/k}-1)/\eta).$
Case $s<0$.
Let $t\in[1,A^{k}]$. As for the case $s>0$, we have $\mathbb{P}\left[|h(\mathbf{X})|\geq\overline{q}t\right]\leq 1/2$. Now since $s<0$, the function $x\mapsto(2^{-s}-1)^{1/s}A^{-1}x^{1/sk}$ is decreasing, and since $t\leq A^{k}$ we have that
$(2^{-s}-1)^{1/s}A^{-1}t^{1/sk}\geq(2^{-s}-1)^{1/s}\geq 1/2,$
which yields that (A.1) holds for all $t\geq 1$ when $s<0$. Now, let $t>0$ and denote $\eta=\eta^{large}_{s,\mathcal{K}}:=\max(1,A/(2^{-s}-1))^{1/s}$. We have that $(2^{-s}-1)^{1/s}A^{-1}t^{1/sk}\leq\eta t^{1/sk}$ and if $t\in(0,1]$ we have that $\eta t^{1/sk}\geq 1$ since $s<0$. Thus, we have the desired result for $s<0$,
$\forall t>0,\quad\mathbb{P}\left[|h(\mathbf{X})|\geq\overline{q}t\right]\leq\eta t^{1/sk}.$
Appendix B Proofs for Section 4
B.1 Proof of Proposition 4.5
Let us start by assuming that $\mathcal{J}(g)=0$. The definition of $\mathcal{J}$ then yields that $\nabla u(\mathbf{X})\in\text{span}\{\nabla g(\mathbf{X})\}$ almost surely. If $\nabla u(\mathbf{X})\neq 0$, then since we are in the one feature case, this implies that $\nabla g(\mathbf{X})\in\text{span}\{\nabla u(\mathbf{X})\}$. Otherwise, we have $\|\nabla u(\mathbf{X})\|_{2}=0$. Hence in both cases, we have
$\|\nabla u(\mathbf{X})\|_{2}^{2}\|\Pi^{\perp}_{\nabla u(\mathbf{X})}\nabla g(\mathbf{X})\|_{2}^{2}=0$
almost surely. As a result $\mathcal{L}_{1}(g)=0$.
Now for the converse statement, let us assume that $\mathcal{L}_{1}(g)=0$. The definition of $\mathcal{L}_{1}$ then yields that $\nabla g(\mathbf{X})\in\text{span}\{\nabla u(\mathbf{X})\}$ almost surely. Then by Assumption 4.3 and since we are in the one feature case, this implies that $\nabla u(\mathbf{X})\in\text{span}\{\nabla g(\mathbf{X})\}$ almost surely, which implies that
$\|\Pi^{\perp}_{\nabla g(\mathbf{X})}\nabla u(\mathbf{X})\|_{2}^{2}=0,$
almost surely. As a result, $\mathcal{J}(g)=0$.
B.2 Proof of Proposition 4.7
Lemma B.1.
Let $a\geq 0$ and $b>0$. Then
$a^{\frac{b}{1+b}}\leq\inf_{x>0}(ax^{-1}+x^{b})\leq 2a^{\frac{b}{1+b}}.$ (B.1)
Proof.
If $a=0$ then the result is straightforward. Now assume that $a>0$ and let
$h:x\mapsto ax^{-1}+x^{b}.$
We have $h\in\mathcal{C}^{\infty}(\mathbb{R}^{*}_{+})$ and $h(x)\rightarrow+\infty$ as $x\rightarrow\{0,+\infty\}$, thus $h$ admits at least 1 global minimum. Now, we have $h^{\prime}(x)=-ax^{-2}+bx^{b-1}$, thus the global minimum of $h$ is at $x=(\frac{a}{b})^{\frac{1}{1+b}}$ and has value $(1+b)(\frac{a}{b})^{\frac{b}{1+b}}$. In other words, we can write
$\inf_{x>0}h(x)=v(b)a^{\frac{b}{1+b}},$
where $v(b)=b\mapsto(1+b)b^{-\frac{b}{1+b}}$. Now we have $v\in\mathcal{C}^{\infty}(\mathbb{R}^{*}_{+})$ and $v(x)\rightarrow 1$ as $x\rightarrow\{0,+\infty\}$. Moreover,
$v^{\prime}(x)=-\frac{\log(x)}{1+x}x^{-\frac{x}{1+x}},$ (B.2)
thus $v$ is strictly increasing on $(0,1)$ and strictly decreasing on $(1,+\infty)$. Thus it holds $21=v(0)\leq v(b)\leq v(1)=$ 2, which yields the desired result. ∎
Let $g\in\mathcal{G}$ and $\alpha>0$. Let the event $E(\alpha):=(\|\nabla g(\mathbf{X})\|_{2}^{2}\leq\alpha)$. From Lemma 4.2 it holds
$\mathcal{J}(g)$ $=\mathbb{E}\left[\|\Pi^{\perp}_{\nabla g(\mathbf{X})}\nabla u(\mathbf{X})\|_{2}^{2}\right]$
$\leq\frac{1}{\alpha}\mathbb{E}\left[\|\nabla u(\mathbf{X})\|_{2}^{2}\|\Pi^{\perp}_{\nabla u(\mathbf{X})}\nabla g(\mathbf{X})\|_{2}^{2}\mathbbm{1}_{\overline{E(\alpha)}}\right]+\mathbb{E}\left[\|\nabla u(\mathbf{X})\|_{2}^{2}\frac{\|\Pi^{\perp}_{\nabla u(\mathbf{X})}\nabla g(\mathbf{X})\|_{2}^{2}}{\|\nabla g(\mathbf{X})\|_{2}^{2}}\mathbbm{1}_{E(\alpha)}\right]$
$\leq\frac{1}{\alpha}\mathcal{L}(g)+\mathbb{E}\left[\|\nabla u(\mathbf{X})\|_{2}^{2}\frac{\|\Pi^{\perp}_{\nabla u(\mathbf{X})}\nabla g(\mathbf{X})\|_{2}^{2}}{\|\nabla g(\mathbf{X})\|_{2}^{2}}.\mathbbm{1}_{E(\alpha)}\right].$
Then using $\|\Pi^{\perp}_{\nabla u(\mathbf{X})}\nabla g(\mathbf{X})\|_{2}^{2}\leq\|\nabla g(\mathbf{X})\|^{2}$ and Holder’s inequality yields
$\mathcal{J}(g)\leq\frac{1}{\alpha}\mathcal{L}(g)+\mathbb{E}\left[\|\nabla u(\mathbf{X})\|_{2}^{2}\mathbbm{1}_{E(\alpha)}\right]\leq\frac{1}{\alpha}\mathcal{L}(g)+\|\|\nabla u(\mathbf{X})\|_{2}^{2}\|_{p_{2}}\mathbb{P}\left[E(\alpha)\right]^{1/p}.$
Now, we will use Proposition 3.3 on $\mathcal{K}:=\{\mathbf{x}\mapsto\|\nabla g(\mathbf{x})\|_{2}^{2}:g\in\mathcal{G}_{1}\}$. Firstly from Assumptions 4.1 and 4.4, $\mathcal{K}$ is a compact set of $\mathcal{C}^{0}\cap L^{1}_{\mu}$ whose elements satisfy (3.2) with constants $k,A<+\infty$. Secondly, by Assumption 4.3 we have $h(\mathbf{X})>0$ almost surely for all $h\in\mathcal{K}$, thus from Proposition 3.2 we have $\underline{q}_{\mathcal{K}}>0$. As a result, we have
$\mathbb{P}\left[E(\alpha)\right]\leq\eta^{small}_{s,\mathcal{K}}(\alpha/\underline{q}_{\mathcal{K}})^{1/k},$
with $\eta^{small}_{s,\mathcal{K}}$ defined in Proposition 3.3. Hence, denoting $\xi:=\big{(}\nu_{u,p_{2}}^{p}\eta^{small}_{s,\mathcal{K}}\underline{q}_{\mathcal{K}}^{-1/k}\big{)}^{1/p}$ it holds
$\mathcal{J}(g)\leq\xi(\alpha^{-1}\frac{\mathcal{L}(g)}{\xi}+\alpha^{1/pk}).$
Finally, the study of $\alpha\mapsto a\alpha^{-1}+\alpha^{b}$ with $a=\mathcal{L}(g)/\xi$ and $b=1/pk$ from Lemma B.1 yields
$\mathcal{J}(g)\leq 2\xi(\frac{\mathcal{L}(g)}{\xi})^{\frac{1}{1+pk}}=2\xi^{\frac{pk}{1+pk}}\mathcal{L}(g)^{\frac{1}{1+pk}}=\gamma_{1}\mathcal{L}(g)^{\frac{1}{1+pk}}$
with $\gamma_{1}$ defined below, recalling that $\eta^{small}_{s,\mathcal{K}}$, $\underline{q}_{\mathcal{K}}$ and $\nu_{u,p_{2}}$ are respectively defined in Proposition 3.3, (3.6) and (4.7),
$\gamma_{1}:=2\big{(}\nu_{u,p_{2}}^{p}\eta^{small}_{s,\mathcal{K}}\underline{q}_{\mathcal{K}}^{-1/k}\big{)}^{\frac{k}{1+pk}}.$ (B.3)
B.3 Proof of Proposition 4.8
Let $g\in\mathcal{G}$ and $\alpha>0$. Let the event $E(\alpha):=(\|\nabla g(\mathbf{X})\|_{2}^{2}\geq\alpha)$. From Lemma 4.2 it holds
$\mathcal{L}(g)$ $\leq\alpha\mathbb{E}\left[\|\Pi^{\perp}_{\nabla g(\mathbf{X})}\nabla u(\mathbf{X})\|_{2}^{2}\mathbbm{1}_{\overline{E(\alpha)}}\right]+\mathbb{E}\left[\|\nabla u(\mathbf{X})\|_{2}^{2}\|\Pi^{\perp}_{\nabla u(\mathbf{X})}\nabla g(\mathbf{X})\|_{2}^{2}\mathbbm{1}_{E(\alpha)}\right]$
$\leq\alpha\mathcal{J}(g)+\mathbb{E}\left[\|\nabla u(\mathbf{X})\|_{2}^{2}\|\Pi^{\perp}_{\nabla u(\mathbf{X})}\nabla g(\mathbf{X})\|_{2}^{2}\mathbbm{1}_{E(\alpha)}\right].$
Then using the fact that $\|\Pi^{\perp}_{\nabla u(\mathbf{X})}\nabla g(\mathbf{X})\|_{2}^{2}\leq\|\nabla g(\mathbf{X})\|_{2}^{2}$ and Holder’s inequality yields
[…]
where $\nu_{u,p_{2}}$ and $\nu_{\mathcal{G}_{1},p_{1}}$ are defined in (4.7). Then, we will use Proposition 3.4 on $\mathcal{K}:=\{\mathbf{x}\mapsto\|\nabla g(\mathbf{x})\|_{2}^{2}:g\in\mathcal{G}_{1}\}$ which is, from Assumptions 4.1 and 4.4, a compact set of $\mathcal{C}^{0}\cap L^{1}_{\mu}$ whose elements satisfy (3.2) with constants $k,A<+\infty$. Note that we will derive results depending if $s\in(0,1/d]$, $s=0$ or if $s<0$. For readability, we will abbreviate $\overline{q}_{\mathcal{K}}$ and $\eta^{large}_{s,\mathcal{K}}$ by $\overline{q}$ and $\eta$.
Case $0<s\leq 1/d$.
Let us start by the simplest case. Indeed if $0<s\leq 1$ then we can actually control $\nu_{g,\infty}=\|\|\nabla g(\mathbf{X})\|_{2}^{2}\|_{\infty}<+\infty$ since $\mathbf{X}$ has compact support. More precisely, taking $\alpha=\overline{q}(1+\eta)^{k}$ yields $\mathbb{P}\left[E(\alpha)\right]=0$, with $\eta$ defined in Proposition 3.4. Hence it holds
$\mathcal{L}(g)\leq\gamma_{2}\mathcal{J}(g)$
with
$\gamma_{2}:=\overline{q}(1+\eta)^{k}.$
Case $s=0$.
Now if $s=0$ we have $\mathbb{P}\left[E(\alpha)\right]\leq\exp(-((\alpha/\overline{q})^{1/k}-1)/\eta)=\exp(1/\eta)\exp(-r\beta)$ where $\beta=\frac{1}{\eta r}(\frac{\alpha}{\overline{q}})^{1/k}$. We can then write
$\mathcal{L}(g)\leq(\eta r)^{k}\overline{q}\mathcal{J}(g)\beta^{k}+\nu_{u,p_{2}}\nu_{\mathcal{G}_{1},p_{1}}\exp(1/\eta r)\exp(-\beta)\leq\kappa(a\beta^{k}+\exp(-\beta))$
with $\kappa:=\nu_{u,p_{2}}\max((\eta r)^{k}\overline{q},\,\nu_{\mathcal{G}_{1},p_{1}}\exp(1/\eta r))$ and $a=\frac{\mathcal{J}(g)}{\nu_{u,p_{2}}}$. Also, by Cauchy-Schwarz inequality we have $a\leq 1$, and one can show that the value of $\beta$ which minimizes the right-hand side behaves like $\log(a^{-1})$ as $a\rightarrow 0$. Hence, we evaluate the right-hand side at $\beta=\log(a^{-1})\geq 0$ to obtain $\mathcal{L}(g)\leq\kappa a(1+\log(a^{-1})^{k})$. In other words,
$\mathcal{L}(g)\leq\gamma_{2}\mathcal{J}(g)(1+\log(\nu_{u,p_{2}}\mathcal{J}(g)^{-1})^{k})\leq\gamma_{2}(1+\log(\nu_{u,p_{2}}\mathcal{J}(g)^{-1}))^{k}$
with
$\gamma_{2}=\nu_{u}\max((\eta r)^{k}\overline{q},\,\nu_{\mathcal{G}_{1},p_{1}}\exp(\frac{1}{\eta r})).$
Case $s<0$.
Finally if $s<0$ then we have $\mathbb{P}\left[E(\alpha)\right]\leq\eta(\alpha/\overline{q})^{1/sk}$, thus
$\mathcal{L}(g)\leq\alpha\mathcal{J}(g)+\kappa\alpha^{1/srk}$
where $\kappa=\nu_{u,p_{2}}\nu_{\mathcal{G}_{1},p_{1}}\eta^{1/r}\overline{q}^{-1/srk}$. Then with $\beta=\alpha^{-1/srk}$ we can write
$\mathcal{L}(g)\leq\mathcal{J}(g)\beta^{-srk}+\kappa\beta^{-1}=\mathcal{J}(g)(a\beta^{-1}+\beta^{b})$
with $a=\frac{\kappa}{\mathcal{J}(g)}\geq 0$ and $b=-srk>0$. Hence applying Lemma B.1 yields
$\mathcal{L}(g)\leq\mathcal{J}(g)2a^{\frac{b}{1+b}}=2\kappa^{\frac{b}{1+b}}\mathcal{J}(g)^{1-\frac{b}{1+b}}=2\kappa^{\frac{-srk}{1-srk}}\mathcal{J}(g)^{\frac{1}{1-srk}}.$
In other words, we have
$\mathcal{L}(g)\leq\gamma_{2}\mathcal{J}(g)^{\frac{1}{1-srk}}$
where
$\gamma_{2}:=2(\nu_{u,p_{2}}\nu_{\mathcal{G}_{1},p_{1}}\eta^{1/r})^{\frac{-srk}{1-srk}}\overline{q}^{\frac{1}{1-srk}}.$
Finally, let us summarize the values for $\gamma_{2}$ depending on the value of $s$, recalling that $\eta=\eta^{large}_{s,\mathcal{K}_{1}}$, $\overline{q}=\overline{q}_{\mathcal{K}_{1}}$ and $\nu$ are respectively defined in Proposition 3.4, (3.6) and (4.7),
$\gamma_{2}:=\left\{\begin{aligned} &\overline{q}(1+\eta)^{k}\quad&\text{if }s\in(0,1/d],\\ &\nu_{u,p_{2}}\max((\eta r)^{k}\overline{q},\,\nu_{\mathcal{G}_{1},p_{1}}\exp(\frac{\eta}{r}))\quad&\text{if }s=0,\\ &2(\nu_{u,p_{2}}\nu_{\mathcal{G}_{1},p_{1}}\eta^{1/r})^{\frac{-srk}{1-srk}}\overline{q}^{\frac{1}{1-srk}}\quad&\text{if }s<0.\\ \end{aligned}\right.$ (B.4)
B.4 Proof of Proposition 4.12
Since $\|\Pi^{\perp}_{\nabla u(\mathbf{X})}\nabla g(\mathbf{X})\|_{2}^{2}=\nabla g(\mathbf{X})^{T}\Pi^{\perp}_{\nabla u(\mathbf{X})}\nabla g(\mathbf{X})$ and $\nabla g(\mathbf{X})=\nabla\Phi(\mathbf{X})G$ it holds
$\mathcal{L}(g)$ $=\mathbb{E}\left[\|\nabla u(\mathbf{X})\|^{2}G^{T}\nabla\Phi(\mathbf{X})^{T}\Pi^{\perp}_{\nabla u(\mathbf{X})}\nabla\Phi(\mathbf{X})G\right]$
$=G^{T}\mathbb{E}\left[\|\nabla u(\mathbf{X})\|^{2}\nabla\Phi(\mathbf{X})^{T}\Pi^{\perp}_{\nabla u(\mathbf{X})}\nabla\Phi(\mathbf{X})\right]G.$
Then since $\Pi^{\perp}_{\nabla u(\mathbf{X})}=I_{d}-\Pi_{\nabla u(\mathbf{X})}$ and $\|\nabla u(\mathbf{X})\|^{2}\Pi_{\nabla u(\mathbf{X})}=\nabla u(\mathbf{X})\nabla u(\mathbf{X})^{T}$ we have
$\mathcal{L}(g)=G^{T}\mathbb{E}\left[\nabla\Phi(\mathbf{X})^{T}\big{(}\|\nabla u(\mathbf{X})\|^{2}I_{d}-\nabla u(\mathbf{X})\nabla u(\mathbf{X})^{T}\big{)}\nabla\Phi(\mathbf{X})\right]G=G^{T}HG.$
The matrix $H\in\mathbb{R}^{K\times K}$ is symmetric as the sum of two symmetric matrices and, since $\mathcal{L}(g)\geq 0$ by definition, it is positive semi-definite.
Appendix C Proofs for Section 5
C.1 Proof of Proposition 5.2
First, recall that, as described in the beginning of Section 5, we have
$\mathcal{J}(g)=\mathbb{E}\left[\|\Pi^{\perp}_{w_{g,j}(\mathbf{X})}v_{g,j}(\mathbf{X})\|_{2}^{2}\right],$
with $w_{g,j}(\mathbf{X})$ and $v_{g,j}(\mathbf{X})$ as defined in (5.1). Moreover, using Lemma 4.2 on the term within the expectation of (5.2) yields
$\mathcal{L}_{m,j}(g)=\mathbb{E}\left[\|w_{g,j}(\mathbf{X})\|_{2}^{2}\|\Pi^{\perp}_{w_{g,j}(\mathbf{X})}v_{g,j}(\mathbf{X})\|_{2}^{2}\right].$
Let us denote $W=\nabla g_{-j}(\mathbf{X})$, and for readability, we will drop the subscripts $g,j$ as well as the dependency in $\mathbf{X}$.
Now, if $\mathcal{J}(g)=0$ then using the first equation yields that almost surely $\|\Pi^{\perp}_{w}v\|_{2}^{2}=0$, and thus using the second equation yields $\mathcal{L}(g)=0$.
Conversely, if $\mathcal{L}(g)=0$, then the second equation implies that almost surely $\|v\|_{2}^{2}=0$ or $\|\Pi^{\perp}_{v}w\|_{2}^{2}=0$. If $\|v\|_{2}^{2}=0$ then $\nabla u(\mathbf{X})\in\text{span}\{(\}W)$, or in other words the gradient of $u$ was already aligned with the gradient of the $m-1$ features $g_{-j}$, thus $\mathcal{J}(g)=0$. Otherwise if $\|\Pi^{\perp}_{v}w\|_{2}^{2}=0$, by Assumption 4.3 we have $w\neq 0$ almost surely, thus it yields $\|\Pi^{\perp}_{w}v\|_{2}^{2}=0$, or in other words
$\Pi^{\perp}_{\Pi^{\perp}_{W}\nabla g_{j}}\Pi^{\perp}_{W}\nabla u=0.$
Finally, we have that $\mathcal{J}(g)=0$ from the fact that
$\Pi^{\perp}_{\Pi^{\perp}_{W}\nabla g_{j}}\Pi^{\perp}_{W}\nabla u=\Pi^{\perp}_{(W|\Pi^{\perp}_{W}\nabla g_{j})}\nabla u=\Pi^{\perp}_{(W|\nabla g_{j})}\nabla u=\Pi^{\perp}_{\nabla g}\nabla u,$
C.2 Proof of Proposition 5.3
For any $\mathbf{x}\in\mathcal{X}$, let us denote $W=\nabla g_{-j}(\mathbf{x})$ and $\mathbf{w}=\nabla g_{j}(\mathbf{x})$. First, by definition of $c$ and $C$ we have that $c\leq\|w\|^{2}_{2}\leq C$. It directly yields that $\|\Pi^{\perp}_{W}\mathbf{w}\|^{2}_{2}\leq C$. Now, let us show that $\|\Pi^{\perp}_{W}\mathbf{w}\|^{2}_{2}\geq c$. We have that for all $\mathbf{y}\in\mathbb{R}^{m-1}$ with $\mathbf{x}=(\mathbf{y}^{T}|-1)^{T}$,
$\|(W|\mathbf{w})\mathbf{x}\|_{2}^{2}\geq c(1+\|\mathbf{y}\|^{2}_{2})\geq c(1+C^{-1}\|W\mathbf{y}\|^{2}_{2}).$
In particular, for $\mathbf{y}=(W^{T}W)^{-1}W^{T}\mathbf{w}$, we have $W\mathbf{y}=\Pi_{W}\mathbf{w}$, thus $(W|\mathbf{w})\mathbf{x}=-\Pi^{\perp}_{W}\mathbf{w}$. Hence, the previous equation reads $\|\Pi^{\perp}_{W}\mathbf{w}\|_{2}^{2}\geq c(1+C^{-1}\|\Pi_{W}\mathbf{w}\|^{2}_{2})$ which yields using $\|\mathbf{w}\|^{2}_{2}\geq c$,
$\|\Pi^{\perp}_{W}\mathbf{w}\|_{2}^{2}\geq c(1+C^{-1}c)-C^{-1}c\|\Pi^{\perp}_{W}\mathbf{w}\|^{2}_{2}.$
This gives us the desired lower bound $\|\Pi^{\perp}_{W}\mathbf{w}\|_{2}^{2}\geq c$. Finally, using (5.2) and (5.3), we obtain the desired result,
$c\mathcal{J}(g)\leq\mathcal{L}_{m,j}(g)\leq C\mathcal{J}(g).$
C.3 Proof of Proposition 5.4
For any $\mathbf{x}\in\mathcal{X}$ let us denote $w(\mathbf{x})=w_{g,j}(\mathbf{x})$ as defined in (5.1) and $W(\mathbf{x})=\nabla g_{-j}(\mathbf{x})$. Note that we dropped the subscript for readability. First, for all $\mathbf{x}\in\mathcal{X}$ we can write
$\|w(\mathbf{x})\|^{2}_{2}=\frac{|W(\mathbf{x})^{T}W(\mathbf{x})|\|\nabla g_{j}(\mathbf{x})\|_{2}^{2}-\nabla g_{j}(\mathbf{x})^{T}W(\mathbf{x})adj(W(\mathbf{x})^{T}W(\mathbf{x}))W(\mathbf{x})^{T}\nabla g_{j}(\mathbf{x})}{|W(\mathbf{x})^{T}W(\mathbf{x})|},$ (C.1)
where $|M|$ denotes the determinant of a matrix $M$, and $adj(M)$ denotes the adjugate matrix of $M$, which is the transposed cofactor matrix of $M$. Now, on $\mathbb{R}^{n\times n}$, we know that $M\mapsto|M|$ and $M\mapsto adj(M)$ are polynomials with total degree at most $n$ and $n-1$ respectively. Hence, $\mathbf{x}\mapsto|W(\mathbf{x})^{T}W(\mathbf{x})|$ and $\mathbf{x}\mapsto adj(W(\mathbf{x})^{T}W(\mathbf{x}))$ are polynomials with total degree at most $2\ell(m-1)$ and $2\ell(m-2)$ respectively. Also, $\mathbf{x}\mapsto\|\nabla g_{j}(\mathbf{x})\|^{2}_{2}$ and $\mathbf{x}\mapsto W^{T}(\mathbf{x})\nabla g_{j}(\mathbf{x})$ are polynomials with total degree at most $2\ell$, hence
$\mathbf{x}\mapsto|W^{T}(\mathbf{x})W(\mathbf{x})|\|\nabla g_{j}(\mathbf{x})\|_{2}^{2},\quad\mathbf{x}\mapsto\nabla g_{j}(\mathbf{x})^{T}W(\mathbf{x})adj(W^{T}(\mathbf{x})W(\mathbf{x}))W^{T}(\mathbf{x})\nabla g_{j}(\mathbf{x})$
are polynomials with total degree at most $2\ell m$. As a result, we can write $\|w(\mathbf{x})\|^{2}_{2}=P_{g,j}(\mathbf{x})/Q_{g,j}(\mathbf{x})$ where
[…]
are multivariate non-negative polynomials with total degrees at most $2\ell m$ and $2\ell(m-1)$ respectively. Now, since the law of $\mathbf{X}$ is $s$-concave with $s>0$, it implies that $\mathcal{X}$ is compact, thus $\sup_{\mathcal{X}}Q_{g,j}<+\infty$ and we can write for all $x\in\mathcal{X}$
$\|w_{g,j}(\mathbf{x})\|^{2}_{2}\geq P_{g,j}(\mathbf{x})(\sup_{\mathcal{X}}Q_{g,j})^{-1}.$
Now, we have two objectives. The first one is to show that $\sup_{\mathcal{X}}Q_{g,j}$ is uniformly bounded for all $g\in\mathcal{G}_{m}$ and $1\leq j\leq m$. The second one is to show that $\mathcal{A}_{m}^{(j)}:=\{P_{g,j},\,g\in\mathcal{G}_{m}\}$ satisfies the assumptions required to apply Propositions 3.2 and 3.3.
Firstly, since $\mathcal{G}_{m}$ is a set of polynomials with bounded total degree, there exists a polynomial $\Phi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{K}$ such that for any $g\in\mathcal{G}_{m}$ we can write $g=G^{T}\Phi$ with $G\in\mathbb{R}^{K\times m}$, which implies
$Q_{g,j}(\mathbf{x})=|G_{-j}^{T}\nabla\Phi(\mathbf{x})^{T}\nabla\Phi(\mathbf{x})G_{-j}|,$
and as a result, for any $1\leq j\leq m$ the function $(G,\mathbf{x})\mapsto Q_{G^{T}\Phi,j}(\mathbf{x})$ is polynomial. Now, since $\mathcal{X}$ and $\mathcal{G}_{m}$ are compact, our first objective is fulfilled,
$\sup_{1\leq j\leq m}\sup_{g\in\mathcal{G}_{m}}\sup_{\mathbf{x}\in\mathcal{X}}|Q_{g,j}(\mathbf{x})|<+\infty.$
Now for the second objective, with the same reasoning, the function $(G,\mathbf{x})\mapsto P_{G^{T}\Phi,j}(\mathbf{x})$ is polynomial, thus continuous, hence since $\mathcal{G}_{m}$ is compact then $\mathcal{A}^{(j)}_{m}$ is also compact in $\mathcal{C}^{0}\cap L^{\infty}$. Finally, we can apply Proposition 3.3 to $\mathcal{A}_{m}^{(j)}$ for any $1\leq j\leq m$, and using the fact that $s>0$ and $A_{P_{g,j}}=4$ we obtain
$\forall\varepsilon>0,\quad\mathbb{P}\left[P_{g,j}(\mathbf{X})\leq\underline{q}_{\mathcal{A}_{m}^{(j)}}\varepsilon\right]\leq 4(1-2^{-s})s^{-1}\varepsilon^{1/2\ell m},$
where $\underline{q}_{\mathcal{A}_{m}^{(j)}}>0$ by Proposition 3.2. As a result, the following small deviation inequality holds for any $g\in\mathcal{G}_{m}$ and $1\leq j\leq m$,
$\mathbb{P}\left[\|w_{g,j}(\mathbf{X})\|^{2}_{2}\leq\underline{q}\varepsilon\right]\leq\eta\varepsilon^{1/2\ell m},$
with
$\underline{q}$ $:=\min_{1\leq j\leq m}\underline{q}_{\mathcal{A}_{m}^{(j)}}>0,$ (C.3)
$\eta$ $=\eta_{s,\mathcal{G}_{m}}^{small}:=4(1-2^{-s})s^{-1}(\sup_{1\leq j\leq m}\sup_{g\in\mathcal{G}_{m}}\sup_{x\in\mathcal{X}}|Q_{g,j}(\mathbf{x})|)^{1/2\ell m}<+\infty.$
C.4 Proof of Proposition 5.5
Firstly, let us state the large deviation inequality satisfied by $\|w_{g,j}(\mathbf{X})\|_{2}^{2}$. Since we have $\|w_{g,j}(\mathbf{X})\|_{2}^{2}\leq\|\nabla g(\mathbf{X})\|_{2}^{2}$ and that $x\mapsto\|\nabla g(\mathbf{x})\|_{2}^{2}$ is a polynomial of total degree at most $2\ell$, we have for all $t>0$,
$\mathbb{P}\left[\|w_{g,j}(\mathbf{X})\|_{2}^{2}\geq\overline{q}t\right]\leq\mathbb{P}\left[\|\nabla g(\mathbf{X})\|_{2}^{2}\geq\overline{q}t\right]\leq(1-(t^{1/2\ell}-1)/\eta)^{1/s}_{+}$
with $\eta=\eta^{large}_{s}=4/(1-2^{-s})$ and $\overline{q}:=\sup_{1\leq j\leq m}\sup_{h\in\mathcal{G}_{m}^{(j)}}Q_{\|\nabla h(\mathbf{X})\|^{2}_{2}}(1/2)$. As a result, we have almost surely that $\|w_{g,j}(\mathbf{X})\|_{2}^{2}\leq\overline{q}(1+\eta)^{2\ell m}$. As a result, we can bound $\mathcal{L}_{m,j}$ by $\mathcal{J}$ as in Proposition 4.8,
$\mathcal{L}_{m,j}(g)\leq\tilde{\gamma}_{2}\mathcal{J}(g).$
with $\tilde{\gamma}_{2}:=\overline{q}(1+\eta)^{2\ell m}$. Secondly, let us control $\mathcal{J}$ using $\mathcal{L}_{m,j}$ thanks to Proposition 5.4. With the exact same reasoning as in Proposition 4.7, with $p=+\infty$ since $\mathbf{X}$ has compact support, we obtain that
$\mathcal{J}(g)\leq\tilde{\gamma}_{1}\mathcal{L}_{m,j}(g)^{\frac{1}{1+2\ell m}}$
with $\tilde{\gamma}_{1}$ of the same form as in (B.3),
$\tilde{\gamma}_{1}:=2(\nu_{u,\infty}\eta_{s,\mathcal{G}_{m}}^{small}\underline{q}^{-1/2\ell m})^{\frac{2\ell m}{1+2\ell m}},$
with $\eta_{s,\mathcal{G}_{m}}^{small}$ and $\underline{q}$ from (C.3). Finally, let $g_{-j}\in\mathcal{G}_{m}^{(-j)}$ and let $g_{j}^{*}\in\mathcal{G}^{(j)}_{m}$ a minimizer of $h\mapsto\mathcal{L}((g_{-j},h))$ over $\mathcal{G}_{m}^{(j)}$. Then for all $h\in\mathcal{G}_{m}^{(j)}$ we have
$\mathcal{J}((g_{-j},g_{j}^{*}))\leq\tilde{\gamma}_{1}\mathcal{L}_{m,j}((g_{-j},g_{j}^{*}))^{\frac{1}{1+2\ell m}}\leq\tilde{\gamma}_{1}\mathcal{L}_{m,j}((g_{-j},h))^{\frac{1}{1+2\ell m}}\leq\tilde{\gamma}_{3}\mathcal{J}((g_{-j},h))^{\frac{1}{1+2\ell m}}$
with $\tilde{\gamma}_{3}=\tilde{\gamma}_{1}\tilde{\gamma}_{2}^{\frac{1}{1+2\ell m}}$.
C.5 Proof of Proposition 5.7
Since $v_{g,j}(\mathbf{X})\in\text{span}\{(\}\nabla g_{-j}(\mathbf{X}))^{\perp}$, the projectors $\Pi^{\perp}_{v_{g,j}(\mathbf{X})}$ and $\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}$ commute. Hence $\|\Pi^{\perp}_{v_{g,j}(\mathbf{X})}w_{g,j}(\mathbf{X})\|_{2}^{2}=\nabla g_{j}(\mathbf{X})^{T}\Pi^{\perp}_{v_{g,j}(\mathbf{X})}\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}\nabla g_{j}(\mathbf{X})$. Also, $\nabla g_{j}(\mathbf{X})=\nabla\Phi(\mathbf{X})G_{j}$, thus
$\mathcal{L}_{m,j}((g_{-j},g_{j}))=G_{j}^{T}\mathbb{E}\left[\|v_{g,j}(\mathbf{X})\|_{2}^{2}\nabla\Phi(\mathbf{X})^{T}\Pi^{\perp}_{v_{g,j}(\mathbf{X})}\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}\nabla\Phi(\mathbf{X})\right]G_{j}.$
Then since $\|v_{g,j}(\mathbf{X})\|_{2}^{2}\Pi^{\perp}_{v_{g,j}(\mathbf{X})}=\|v_{g,j}(\mathbf{X})\|_{2}^{2}I_{d}-v_{g,j}(\mathbf{X})v_{g,j}(\mathbf{X})^{T}$ we have
$\mathcal{L}((g_{-j},g_{j}))=G_{j}^{T}\mathbb{E}\left[\nabla\Phi(\mathbf{X})^{T}\big{(}\|v_{g,j}(\mathbf{X})\|^{2}\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}-v_{g,j}(\mathbf{X})v_{g,j}(\mathbf{X})^{T}\big{)}\nabla\Phi(\mathbf{X})\right]G_{j},$
in other words $\mathcal{L}((g_{-j},g_{j}))=G_{j}^{T}H_{m,j}G_{j}$ where $H_{m,j}\in\mathbb{R}^{K\times K}$ as defined in (5.11), is symmetric, as the sum of two symmetric matrices, and since $\mathcal{L}_{m,j}(g)\geq 0$ by definition, it is positive semi-definite.

### v2

@@
 The constant $C(\mathbf{X}|\mathcal{G}_{m})$ is defined via probabilistic Poincaré constants associated to the conditional random variables $\mathbf{X}|G^{T}\mathbf{X}=\mathbf{z}$ for $\mathbf{z}\in\mathbb{R}^{m}$.
-The fact that the support of the latter lies in the $(d-m)$-dimensional affine space $\mathbf{z}+\text{span}\{(\}G)^{\perp}$ allows to derive bounds on $C(\mathbf{X}|\mathcal{G}_{m})$ for several classical probability measures [23].
+The fact that the support of the latter lies in the $(d-m)$-dimensional affine space $\mathbf{z}+\text{span}\{G\}^{\perp}$ allows to derive bounds on $C(\mathbf{X}|\mathcal{G}_{m})$ for several classical probability measures [24].
 The other main advantage of this method is that, while minimizing $\mathcal{E}$ on $\mathbb{R}^{d\times m}$ is a challenging problem, minimizing $\mathcal{J}$ on $\mathbb{R}^{d\times m}$ is rather straightforward.
 Indeed, it is a quadratic functional, whose minimum is the sum of the $d-m$ smallest eigenvalues of the matrix $\mathbb{E}\left[\nabla u(\mathbf{X})\nabla u(\mathbf{X})^{T}\right]\in\mathbb{R}^{d\times d}$, and the minimizer is a matrix whose columns are the eigenvectors associated with the $m$ largest eigenvalues of that matrix.
-Although linear dimension reduction has several theoretical and practical advantages, the main limitation is that it may not efficiently detect important nonlinear features, as it is easy to see by considering $u()\mathbf{x})=\mapsto h(\|\mathbf{x}\|^{2}_{2})$ for some $h\in\mathcal{C}^{1}$.
+Although linear dimension reduction has several theoretical and practical advantages, the main limitation is that it may not efficiently detect important nonlinear features, as it is easy to see by considering $u(\mathbf{x})=h(\|\mathbf{x}\|^{2}_{2})$ for some $h\in\mathcal{C}^{1}$.
 In this regard, there are insights for developing nonlinear dimension reduction methods, where the model class $\mathcal{G}_{m}$ contains nonlinear functions.
@@
 In particular in [3] it was shown that (1.2) still holds, under important assumptions, with
-$\mathcal{J}(g)=\mathbb{E}\left[\|\Pi^{\perp}_{\text{span}\{\nabla\}g(\mathbf{X})}\nabla u(\mathbf{X})\|^{2}_{2}\right]=\mathbb{E}\left[\|\nabla u(\mathbf{X})\|^{2}_{2}\right]-\mathbb{E}\left[\|\Pi_{\text{span}\{\nabla\}g(\mathbf{X})}\nabla u(\mathbf{X})\|^{2}_{2}\right].$ (1.3)
+$\mathcal{J}(g)=\mathbb{E}\left[\|\Pi^{\perp}_{\text{span}\{\nabla g(\mathbf{X})\}}\nabla u(\mathbf{X})\|^{2}_{2}\right]=\mathbb{E}\left[\|\nabla u(\mathbf{X})\|^{2}_{2}\right]-\mathbb{E}\left[\|\Pi_{\text{span}\{\nabla g(\mathbf{X})\}}\nabla u(\mathbf{X})\|^{2}_{2}\right].$ (1.3)
 The constant $C(\mathbf{X}|\mathcal{G}_{m})$ is here defined via probabilistic Poincaré inequalities on manifolds, associated to the conditional random variables $\mathbf{X}|g(\mathbf{X})=\mathbf{z}$ for $\mathbf{z}\in\mathbb{R}^{m}$.
 Indeed, the support of the latter lies in $g^{-1}(\{\mathbf{z}\})$, which is no longer an affine space but rather a $(d-m)$-dimensional smooth submanifold of $\mathbb{R}^{d}$, under some assumptions on $g$.
-Note that $\Pi^{\perp}_{\text{span}\{\nabla\}g(\mathbf{X})}\nabla u(\mathbf{X})$ is actually the Riemanian gradient of $u_{|g^{-1}(\{g(\mathbf{X})\})}$ at point $\mathbf{X}$, associated to the euclidean metric in $\mathbb{R}^{d}$.
+Note that $\Pi^{\perp}_{\text{span}\{\nabla g(\mathbf{X})\}}\nabla u(\mathbf{X})$ is actually the Riemanian gradient of $u_{|g^{-1}(\{g(\mathbf{X})\})}$ at point $\mathbf{X}$, associated to the euclidean metric in $\mathbb{R}^{d}$.
 This nonlinear setting comes with two main issues.
@@
 The second issue is that optimizing $\mathcal{J}$ over a class $\mathcal{G}_{m}$ of nonlinear functions can be a challenging task, although first and second order optimization algorithms are available.
-This is mainly due to the nonlinearity of the mapping $g\mapsto\Pi^{\perp}_{\text{span}\{\nabla\}g(\mathbf{x})}$.
+This is mainly due to the nonlinearity of the mapping $g\mapsto\Pi^{\perp}_{\text{span}\{\nabla g(\mathbf{x})\}}$.
 This second issue is the motivation for the present work, in which we derive new upper bounds that yield to straightforward optimization procedures.
@@
 $\nabla u(\mathbf{x})\in\text{span}\{\nabla g(\mathbf{x})\}:=\text{span}\{\nabla g_{i}(\mathbf{x})\}_{1\leq i\leq m}\subset\mathbb{R}^{d}.$
-Hence there is an incentive to build $g$ such that $\nabla u(\mathbf{x})$ and $\text{span}\{\nabla\}g(\mathbf{x})$ are as aligned as possible.
+Hence there is an incentive to build $g$ such that $\nabla u(\mathbf{x})$ and $\text{span}\{\nabla{g(\mathbf{x})}\}$ are as aligned as possible.
 The authors in [3] propose to minimize the following cost function,
 $\mathcal{J}(g):=\mathbb{E}\left[\|\nabla u(\mathbf{X})\|^{2}\right]-\mathbb{E}\left[\|\Pi_{\nabla g(\mathbf{X})}\nabla u(\mathbf{X})\|^{2}\right],$ (2.1)
-where $\Pi_{\nabla g(\mathbf{x})}:=\Pi_{\text{span}\{\nabla\}g(\mathbf{x})}\in\mathbb{R}^{d\times d}$ denotes the orthogonal projector from $\mathbb{R}^{d}$ onto $\text{span}\{\nabla\}g(\mathbf{x})$.
+where $\Pi_{\nabla g(\mathbf{x})}:=\Pi_{\text{span}\{\nabla g(\mathbf{x})\}}\in\mathbb{R}^{d\times d}$ denotes the orthogonal projector from $\mathbb{R}^{d}$ onto $\text{span}\{\nabla g(\mathbf{x})\}$.
 The cost function $\mathcal{J}(g)$ naturally appears when using Poincaré inequalities and provides certified error bounds for $g\in\mathcal{G}_{m}$ [28].
@@
 In this case, for any $\mathbf{x}\in\mathcal{M}_{\mathbf{z}}$, it holds
-$T_{x}\mathcal{M}_{\mathbf{z}}=\text{ker}\nabla g(\mathbf{x})^{T}=(\mathrm{span}\nabla g(\mathbf{x}))^{\perp}$
+$T_{\mathbf{x}}\mathcal{M}_{\mathbf{z}}=\text{ker}\{\nabla g(\mathbf{x})^{T}\}=\text{span}\{\nabla g(\mathbf{x})\}^{\perp}$
 and thus
@@
 It is important to note that $\mathcal{J}$ satisfies some invariance properties.
-Indeed, for any diffeomorphism $\psi:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$, by the chain rule and invertibility of $\nabla\psi(\mathbf{x})$, it holds that $\text{span}\{\nabla\}(\psi\circ g(\mathbf{x}))=\text{span}\{\nabla\}g(\mathbf{x})$ and thus
+Indeed, for any diffeomorphism $\psi:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$, by the chain rule and invertibility of $\nabla\psi(\mathbf{x})$, it holds that $\text{span}\{\nabla(\psi\circ g(\mathbf{x}))\}=\text{span}\{\nabla g(\mathbf{x})\}$ and thus
 $\mathcal{J}(\psi\circ g)=\mathcal{J}(g).$ (2.5)
@@
 In this section, we propose an extension of the surrogate we introduced in (4.1) to the general case $m\geq 1$.
-The main problem when trying to do so is that for $m>1$, $\text{span}\{\nabla\}u(\mathbf{X})$ and $\text{span}\{\nabla\}g(\mathbf{X})$ are no longer of the same dimension, thus we can no longer expect to relate $\|\Pi^{\perp}_{\nabla g(\mathbf{X})}\nabla u(\mathbf{x})\|_{2}$ with $\|\Pi^{\perp}_{\nabla u(\mathbf{x})}\nabla g(\mathbf{X})\|_{F}$ in a similar way as in Lemma 4.2.
+The main problem when trying to do so is that for $m>1$, $\text{span}\{\nabla u(\mathbf{X})\}$ and $\text{span}\{\nabla g(\mathbf{X})\}$ are no longer of the same dimension, thus we can no longer expect to relate $\|\Pi^{\perp}_{\nabla g(\mathbf{X})}\nabla u(\mathbf{x})\|_{2}$ with $\|\Pi^{\perp}_{\nabla u(\mathbf{x})}\nabla g(\mathbf{X})\|_{F}$ in a similar way as in Lemma 4.2.
 For this reason, the extension we propose relies on learning one feature at a time, in a coordinate descent or greedy approach.
@@
 Proof.
-The projectors $\Pi^{\perp}_{v_{g,j}(\mathbf{X})}$ and $\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}$ commute since $v_{g,j}(\mathbf{X})\in\text{span}\{(\}\nabla g_{-j}(\mathbf{X}))^{\perp}$.
+The projectors $\Pi^{\perp}_{v_{g,j}(\mathbf{X})}$ and $\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}$ commute since $v_{g,j}(\mathbf{X})\in\text{span}\{\nabla g_{-j}(\mathbf{X})\}^{\perp}$.
 The definition of $\mathcal{L}_{m,j}$ in (5.2) then yields the result.
@@
 One may also consider optimizing both in one global cross-validation procedure, although it would be much more costly.
-Implementation.
+Implementation and data statement.
 The cross-validation procedures as well as the Kernel ridge regression rely on the library sklearn [21].
@@
 The orthonormal polynomials feature maps rely on the python library tensap [19].
-Open-source implementation of the approach proposed in this work, as well as tutorials, are available at https://github.com/alexandre-pasco/tensap/tree/poincare_learning.
+The code underlying this article is freely available in [20] at https://doi.org/10.5281/zenodo.15430309.
 6.2 The case of a single feature
@@
 Conversely, if $\mathcal{L}(g)=0$, then the second equation implies that almost surely $\|v\|_{2}^{2}=0$ or $\|\Pi^{\perp}_{v}w\|_{2}^{2}=0$.
-If $\|v\|_{2}^{2}=0$ then $\nabla u(\mathbf{X})\in\text{span}\{(\}W)$, or in other words the gradient of $u$ was already aligned with the gradient of the $m-1$ features $g_{-j}$, thus $\mathcal{J}(g)=0$.
+If $\|v\|_{2}^{2}=0$ then $\nabla u(\mathbf{X})\in\text{span}\{W\}$, or in other words the gradient of $u$ was already aligned with the gradient of the $m-1$ features $g_{-j}$, thus $\mathcal{J}(g)=0$.
 Otherwise if $\|\Pi^{\perp}_{v}w\|_{2}^{2}=0$, by Assumption 4.3 we have $w\neq 0$ almost surely, thus it yields $\|\Pi^{\perp}_{w}v\|_{2}^{2}=0$, or in other words
@@
 C.5 Proof of Proposition 5.7
-Since $v_{g,j}(\mathbf{X})\in\text{span}\{(\}\nabla g_{-j}(\mathbf{X}))^{\perp}$, the projectors $\Pi^{\perp}_{v_{g,j}(\mathbf{X})}$ and $\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}$ commute.
+Since $v_{g,j}(\mathbf{X})\in\text{span}\{\nabla g_{-j}(\mathbf{X})\}^{\perp}$, the projectors $\Pi^{\perp}_{v_{g,j}(\mathbf{X})}$ and $\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}$ commute.
 Hence $\|\Pi^{\perp}_{v_{g,j}(\mathbf{X})}w_{g,j}(\mathbf{X})\|_{2}^{2}=\nabla g_{j}(\mathbf{X})^{T}\Pi^{\perp}_{v_{g,j}(\mathbf{X})}\Pi^{\perp}_{\nabla g_{-j}(\mathbf{X})}\nabla g_{j}(\mathbf{X})$.